### v1

# ATLAS: Adaptive Trading with LLM AgentS Through Dynamic Prompt Optimization and Multi-Agent Coordination

Abstract
Large language models show promise for financial decision-making, yet deploying them as autonomous trading agents raises fundamental challenges: how to adapt instructions when rewards arrive late and obscured by market noise, how to synthesize heterogeneous information streams into coherent decisions, and how to bridge the gap between model outputs and executable market actions. We present ATLAS (Adaptive Trading with LLM AgentS), a unified multi-agent framework that integrates structured information from markets, news, and corporate fundamentals to support robust trading decisions. Within ATLAS, the central trading agent operates in an order-aware action space, ensuring that outputs correspond to executable market orders rather than abstract signals. The agent can incorporate feedback while trading using Adaptive-OPRO, a novel prompt-optimization technique that dynamically adapts the prompt by incorporating real-time, stochastic feedback, leading to increasing performance over time. Across regime-specific equity studies and multiple LLM families, Adaptive-OPRO consistently outperforms fixed prompts, while reflection-based feedback fails to provide systematic gains [^1] Code will be available upon publication..
ATLAS: Adaptive Trading with LLM AgentS
Through Dynamic Prompt Optimization and Multi-Agent Coordination
Charidimos Papadakis, Angeliki Dimitriou, Giorgos Filandrianos, Maria Lymperaiou, Konstantinos Thomas, Giorgos Stamou School of Electrical and Computer Engineering, AILS Laboratory National Technical University of Athens harrypapadakis02@gmail.com, {angelikidim, geofila, marialymp, kthomas}@ails.ece.ntua.gr, gstam@cs.ntua.gr
1 Introduction
Financial markets represent one of humanity’s most complex decision-making environments, requiring synthesis of vast information from technical indicators and fundamental analysis to breaking news and market sentiment. Large Language Models (LLMs) introduce new possibilities for financial decision-making through their ability to process diverse data sources and reason over complex scenarios.
From the model’s perspective, financial trading serves as an ideal testbed due to its combination of unambiguous metrics, sequential complexity, multimodal reasoning requirements, and inherent stochasticity. Unlike synthetic benchmarks, markets provide extensive historical data without simulation bias and reward genuine understanding over pattern memorization. LLMs can therefore be tasked to make decisions under uncertainty, revealing capabilities in complex reasoning He et al. (2025), market understanding Li et al. (2025), and high-risk decision-making Hung et al. (2023).
Turning these capabilities into reliable systems raises practical questions: i) how a model adapts with delayed, noisy rewards, ii) how diverse signals form coherent guidance, and iii) how decisions are expressed to align with executable market actions. Recent progress touches parts of this space, yet adaptation under delayed feedback and action space expressivity remain under-specified.
We introduce ATLAS (Adaptive Trading with LLM AgentS) to address these points. ATLAS centers on Adaptive-OPRO, a prompt optimization mechanism for sequential settings, and an order-aware decision layer that ties model outputs to execution. Structured analyses of market data, news, and fundamentals provide consistent inputs to the decision policy, while optimized instructions determine how these inputs are synthesized over time. Our experiments show that different LLMs develop distinct trading behaviors and that the combination of adaptive prompting with coordinated inputs improves decision quality over fixed prompts. Crucially, we find that reflection-based reasoning, although often assumed to be beneficial, can degrade performance in well-tuned systems, and that single-run evaluations can conceal substantial variance and lead to unstable conclusions.
These findings outline practical principles for deploying LLMs in high-stakes settings with sequential decisions under real-world uncertainty. In parallel, ATLAS improves explainability of agent decisions through explicit order specifications, structured analyst inputs, and transparent prompt evolution, supporting collaboration with financial experts and more reliable deployment.
2 Related Work
LLM Agents in Financial Markets
Recent work explores several LLM-based trading agents, from sentiment-driven pipelines Kirtac and Germano (2024) to coordinated, multi-component systems Zhou et al. (2025); Yang et al. (2025); Liu et al. (2023). Examples include CryptoTrade, which integrates on/off-chain signals with a reflection component Li et al. (2024), and TradingAgents, which organizes fundamental, sentiment, and technical analysts with debate mechanisms Xiao et al. (2025). Memory-centric designs such as FinMem emphasize persistent, task-specific recall Yu et al. (2023), while FINCON introduces conceptual verbal reinforcement to shape multi-agent collaboration Yu et al. (2024). Other directions incorporate learning signals Xiong et al. (2025) or mixture-of-experts routing Ding et al. (2025), and focus on document-centric analysis such as filings and earnings calls Fatouros et al. (2025). Despite advances, issues remain: manually crafted prompts with no systematic adaptation under delayed, noisy feedback, and simplified action spaces that collapse execution into directional scores. We therefore study order-level outputs (type, size, timing, price) in a simulator built for such interfaces Papadakis et al. (2025), and adopt multi-run reporting to address stochastic variability Song et al. (2025); Atil et al. (2025).
Prompt Engineering and Optimization
Prompt optimization enhances LLM performance beyond manual tuning. Optimization by PROmpting (OPRO) treats the model as a meta-optimizer over instruction text and has shown gains on single-turn tasks with immediate feedback Yang et al. (2024). Extensions explore evolutionary search and reinforcement-style updates Guo et al. (2025); Do et al. (2024); Austin and Chartock (2024). These settings typically assume fast, unambiguous scoring and independent instances. In contrast, trading provides deferred, noisy reward signals and sequentially coupled decisions. Adaptive-OPRO adapts prompt optimization to this regime by using rolling evaluation windows and by separating static instructions from dynamic run-time content, allowing stability where consistency matters and controlled evolution where change is beneficial.
3 ATLAS Framework
ATLAS comprises three main components: (i) a Market Intelligence Pipeline, which consists of specialized agents that prepare market, news, and fundamental inputs for downstream decisions; (ii) a Decision & Execution Layer centered on a Central Trading Agent that generates and executes orders; and (iii) a feedback mechanism that collects post-execution signals and feeds them back for continuous adaptation. Within the feedback mechanism we introduce Adaptive-OPRO, an extension of the OPRO framework that dynamically edits the Central Trading Agent’s instruction prompt based on real-time, stochastic market feedback. Figure 1 provides an overview of the ATLAS framework.
3.1 Market Intelligence Pipeline
ATLAS separates information preparation from decision-making. Thus, the Market Intelligence Pipeline consists of three distinct agents, each with a specialized analyst role. Market Analyst produces multi-timescale summaries from price and volume in varying time scales (2 years, 6 months, and 3 months of history with monthly, weekly, and daily candlesticks, respectively). Within each window it computes standard indicators (e.g., moving averages, momentum, volatility bands, support/resistance) and refreshes daily, providing a consistent, noise-filtered description rather than trading signals (details in App. B). News Analyst aggregates relevant articles into structured fields (Sentiment Assessment, Key Developments, Market Relevance, Source Analysis) with optional full-text retrieval to move beyond headlines (details in App.  C.1). Fundamental Analyst extracts material changes from periodic reports and corporate events, activating infrequently to mirror reporting cycles and provide medium- to long-horizon context (details in App.  C.2).
Figure 1: ATLAS Framework Overview. The Central Trading Agent submits orders to the Trading Execution Engine via prompts shaped by three specialized analysts and the proposed Adaptive-OPRO optimization technique.
3.2 Decision & Execution Layer
The Decision & Execution Layer determines trading actions (e.g., buying or selling a stock), executes these orders, and receives corresponding market feedback. The main decision-making component within this layer is the Central Trading Agent. Central Trading Agent consumes the structured inputs and current portfolio and emits orders that specify type (market, limit, stop), size, timing, and price levels. Orders are executed in StockSim Papadakis et al. (2025), which enforces core trading semantics and returns fills, positions, and cash for the next step. Order-level decisions clarify intent and link analytical quality to execution choices.
3.3 Adaptive-OPRO
Adaptive-OPRO updates the prompt of the Central Trading Agent using realized outcomes, targeting sequential decisions with late, noisy reward feedback. It edits only the agent’s instruction text while maintaining analyst texts, addressing three needs: credit assignment under delay (updates reflect decision sequences rather than single steps), stability under noise (avoiding short-lived fluctuations), and edit locality (preserving the run-time interface that injects analyst outputs and portfolio state).
Template separation.
The Central Trading Agent prompt is a template with two parts: static instructions (decision criteria, risk preferences, output schema) and dynamic run-time content (current Market/News/Fundamental summaries, portfolio, recent orders). Adaptive-OPRO edits only static instructions; placeholders for dynamic fields are preserved, preventing overfitting to transient data and keeping the execution path intact (Appendix G).
Windowed scoring and choice of signal.
Feedback is aggregated over windows of $K=5$ decision steps. For each window, cumulative return on investment (ROI) is computed and mapped to a score $s\in[0,100]$ following the OPRO scoring convention (Eq 1) Yang et al. (2024), so that $20\%\!\mapsto\!0$, $0\%\!\mapsto\!50$, $20\%\!\mapsto\!100$.
$s=\mathrm{clip}_{[0,100]}\big(50+250\cdot\mathrm{ROI}\big)$ (1)
The choice of $K$ and this linear scaling is carefully selected for our evaluation period and can be adjusted for other horizons or volatility regimes without change. ROI is directly aligned with the objective (capital growth) and is deterministically computed from portfolio value; when windowed and clipped it provides a low-variance, delay-aware signal that reflects the cumulative effect of a sequence of decisions rather than any single outcome.
Meta-prompted update.
At each window boundary, Adaptive-OPRO builds a prompt evolution history containing prior instruction variants and their scores. A meta-prompt presents this history to an optimizer LLM and requests four outputs: (i) a brief performance analysis that identifies likely bottlenecks, (ii) a proposed instruction update, (iii) a concise list of key improvements, and (iv) an expected impact summary. This analysis makes the rationale for edits explicit and ties changes to observed behavior, clearly showing how the policy evolves. The candidate replaces the current instruction block if the template’s placeholder set is unchanged, ensuring compatibility with the run-time injector.
Model Prompting ROI (%) $\uparrow$ SR $\uparrow$ DD (%) $\downarrow$ Win Rate (%) $\uparrow$ Num Trades
Non-LLM-Based Strategies
Buy & Hold N/A -8.59 -0.071 20.45 0.00 1
MACD N/A 6.50 0.131 6.86 0.00 1
SMA N/A 6.91 0.177 3.56 50.00 4
SLMA N/A -1.87 -0.078 6.89 0.00 1
Bollinger Bands N/A 0.00 0.000 0.00 0.00 0
LLM-Based Strategies - ATLAS
LLaMA 3.3-70B Baseline -9.19± 1.54 -0.091± 0.021 16.90± 0.82 30.28± 11.87 22.67± 8.39
Reflection -8.44± 1.58 -0.087± 0.025 16.36± 0.31 44.69± 13.25 27.67± 1.15
Adaptive-OPRO -6.16± 2.08 -0.066± 0.004 14.05± 3.33 54.36± 12.44 28.33± 3.21
Qwen3-235B Baseline -1.78± 3.86 -0.006± 0.039 13.09± 1.88 36.51± 17.55 13.00± 4.00
Reflection -5.76± 2.97 -0.049± 0.033 14.18± 1.91 25.00± 0.00 8.67± 0.58
Adaptive-OPRO 1.33± 1.91 0.025± 0.019 11.41± 0.06 50.00± 0.00 9.00± 0.00
Qwen3-32B Baseline -10.62± 3.54 -0.087± 0.031 16.72± 2.75 30.00± 10.00 25.33± 1.53
Reflection -7.76± 0.90 -0.065± 0.002 16.47± 3.44 28.72± 25.06 31.67± 2.31
Adaptive-OPRO -3.48± 2.19 -0.022± 0.021 15.52± 0.68 43.45± 6.27 28.67± 1.53
Claude Sonnet 4
Baseline -7.26± 2.99 -0.066± 0.030 17.59± 1.55 31.19± 7.84 13.00± 4.36
Reflection -5.69± 1.82 -0.058± 0.013 15.12± 3.26 46.67± 5.77 12.67± 2.08
Adaptive-OPRO 0.35± 1.78 0.008± 0.018 14.76± 2.87 43.45± 6.27 15.00± 2.00
Claude Sonnet 4
w/ Thinking
Baseline -4.46± 4.76 -0.043± 0.048 14.32± 4.12 11.11± 19.24 14.00± 2.65
Reflection -8.60± 0.59 -0.078± 0.004 19.45± 1.65 14.29± 24.75 11.67± 2.08
Adaptive-OPRO -0.73± 3.82 -0.004± 0.038 12.94± 2.32 43.89± 21.11 17.00± 5.00
GPT-o4-mini Baseline -1.30± 1.71 -0.017± 0.017 9.68± 3.12 29.17± 11.02 15.33± 3.06
Reflection -2.52± 4.03 -0.039± 0.045 9.82± 3.43 51.28± 5.06 20.33± 3.06
Adaptive-OPRO 9.06± 0.73 0.094± 0.008 11.48± 0.00 65.28± 16.84 17.33± 5.86
GPT-o3 Baseline -6.11± 3.42 -0.080± 0.029 11.58± 3.09 42.59± 8.49 18.67± 3.21
Reflection -4.60± 3.40 -0.053± 0.044 12.11± 1.27 46.03± 16.88 18.33± 2.52
Adaptive-OPRO 9.02± 3.28 0.146± 0.048 5.33± 0.14 72.81± 17.27 19.67± 4.16
Table 1: Performance comparison between non-LLM-based and LLM-based approaches using ATLAS in volatile, declining market conditions (LLY, healthcare sector). Bold values indicate the best per model.
4 Experiments
Our study examines ATLAS along three axes: (1) Adaptation – whether sequential prompt optimization via Adaptive-OPRO improves over strong fixed prompts and over analytical reflection when feedback is delayed and noisy; (2) Component attribution – the contribution of structured inputs (Market Analyst, News Analyst, Fundamental Analyst) under different regimes; (3) Model capabilities – how backbone LLMs perform as both decision policies and prompt optimizers under Adaptive-OPRO, assessed by return and risk-adjusted performance, robustness across runs, and their ability to propose instruction updates that yield sustained improvements over windows.
4.1 Experimental Design
Assets and timeperiod.
We evaluate listed equities in three regime-specific windows that emphasize distinct conditions: bearish-volatile (LLY), sideways (XOM), and bullish (NVDA). Each window spans two months (Apr 28-Jun 28, 2025) with a daily decision interval: the agent may act once per trading day. This horizon is chosen to (i) capture multiple decision cycles without regime mixing, so adaptation reflects outcomes rather than macro shifts, and (ii) preserve complete conversation history (analyst summaries, orders, prompt-evolution logs) within the context limits of all backbones, enabling fair, auditable runs across models and ablations.
To avoid outcome-driven selection, assets were identified in advance using simple, transparent criteria (liquidity, sector diversity, representative regime behavior). ATLAS is horizon- and asset-agnostic; the same protocol applies to alternative windows or instruments, with the scoring window and scale chosen to match the volatility profile.
Models.
We evaluate seven backbones spanning families, sizes, and reasoning modes: GPT-o3, GPT-o4-mini, Claude Sonnet 4 with and without thinking, LLaMA 3.3-70B, Qwen3-235B, and Qwen3-32B. Within a run, the same backbone powers all ATLAS components to avoid cross-model effects. This mix lets us examine how capability and architecture shape sequential decision making, instruction adherence, and stability, and it tests whether Adaptive-OPRO transfers across model families without per-model tuning.
Prompting strategies.
We compare three strategies for the Central Trading Agent: Baseline – a strong fixed prompt produced via iterative expert prompt engineering; Reflection – a weekly reflection mechanism adapted from Li et al. (2024) that generates analytical feedback the agent must interpret; Adaptive-OPRO – our sequential prompt optimization with windowed scoring and template separation (Section 3.3). Our objective is to isolate the adaptation mechanism under identical data and execution semantics, so we use a single, transparent experimental setup rather than re-creating full external agent stacks, which differ in action-space definitions, state handling, and execution interfaces and would confound comparisons. Reflection is included as a widely used, portable form of sequential feedback, providing a focused and fair point of contrast to Adaptive-OPRO and the fixed baseline.
Non-LLM baselines.
Following Li et al. (2024), we include five widely used quantitative strategies to contextualize results: Buy & Hold, MACD Wang and Kim (2018), SMA Gencay (1996), SLMA Wang and Kim (2018), and Bollinger Bands Day et al. (2023). For window-based methods, we test multiple window lengths per regime and report a strong, representative configuration for each strategy (e.g., 10-day SMA; 10/30-day SLMA). Full specifications appear in Appendix D.7.
Execution environment.
Agents interact with StockSim Papadakis et al. (2025) through an order-level interface. The simulator enforces core trading semantics (submitting orders with price levels and sizes, updating positions and cash, and reporting fills) and produces a complete audit trail of orders, executions, and portfolio state. This setup moves beyond generic directional views and evaluates whether a policy can translate analysis into concrete, executable actions. We intentionally abstract fine-grained market microstructure to keep runs comparable and to isolate the decision policy. The objective of our evaluation is order-aware reasoning and coherent execution under standard semantics, not microstructure modeling. Fixing these mechanics makes results deterministic and auditable across models and ablations.
Evaluation Metrics.
We employ 5 metrics capturing different aspects of trading performance:
Return on Investment (ROI): Total percentage return calculated as: $\frac{\text{final value}-\text{initial value}}{\text{initial value}}\times 100$, where portfolio values include both cash holdings and the current market value of all stocks owned.
Sharpe Ratio (SR): Risk-adjusted return metric calculated as: $\frac{\mu-r_{f}}{\sigma}$, where $\mu$ is mean daily return, $\sigma$ is daily return standard deviation, and $r_{f}$ is the risk-free rate (set to 0 following Li et al. (2024)). Higher values indicate better returns per unit of risk.
Maximum Drawdown (DD): The worst peak-to-trough decline in portfolio value: $\max_{t\in[0,T]}\left(\max_{s\in[0,t]}V_{s}-V_{t}\right)/\max_{s\in[0,t]}V_{s}$, where $V_{t}$ is portfolio value at time $t$. This measures the largest loss from any historical high point, capturing the strategy’s downside risk and stress tolerance.
Win Rate: Percentage of closed (i.e., completed) trades that are profitable:
[…]
Number of Trades: Total trading frequency over the evaluation period. Higher frequencies indicate active, opportunistic short-term strategies, while lower frequencies suggest patient, conviction-driven approaches. Additional metrics, results, and analyses are reported in Appendix E.
Protocol and statistics.
To address the stochasticity of LLM outputs, each configuration is run three times and we report results as mean $\pm$ standard deviation. This multiple-run protocol captures variability and helps distinguish genuine capability differences from random fluctuations. Together with the reported distributional and execution diagnostics, it surfaces behavioral aspects that single-run summaries may miss.
Model Prompting XOM NVDA
ROI (%) $\uparrow$ SR $\uparrow$ DD (%) $\downarrow$ ROI (%) $\uparrow$ SR $\uparrow$ DD (%) $\downarrow$
Non-LLM-Based Strategies
Buy & Hold N/A 1.14 0.013 6.97 41.30 0.409 3.16
MACD N/A -0.26 -0.019 5.90 -0.62 -0.343 0.62
SMA N/A -1.02 -0.019 5.75 14.02 0.242 2.93
SLMA N/A -2.08 -0.066 5.53 36.77 0.386 3.12
Bollinger Bands N/A 0.00 0.000 0.00 0.00 0.000 0.00
LLM Based-Strategies - ATLAS
LLaMA 3.3-70B Baseline -0.42± 2.06 -0.024± 0.051 5.56± 1.08 37.86± 12.31 0.388± 0.096 3.46± 0.63
Reflection -2.61± 0.77 -0.083± 0.014 6.38± 0.72 40.40± 1.43 0.422± 0.023 2.96± 0.34
Adaptive-OPRO -1.10± 0.44 -0.045± 0.012 5.15± 0.71 42.07± 1.85 0.418± 0.016 3.15± 0.02
Qwen3-235B Baseline -2.43± 0.68 -0.044± 0.014 5.72± 0.15 43.91± 2.31 0.416± 0.001 3.34± 0.16
Reflection -2.02± 1.44 -0.037± 0.034 6.26± 1.77 34.08± 12.30 0.374± 0.075 2.98± 0.30
Adaptive-OPRO 0.27± 1.83 0.011± 0.037 7.20± 2.09 41.25± 0.00 0.418± 0.000 3.16± 0.00
Qwen3-32B Baseline -9.14± 1.02 -0.204± 0.023 9.82± 0.90 35.75± 5.35 0.477± 0.060 2.86± 0.30
Reflection -7.96± 3.11 -0.162± 0.060 9.05± 2.90 41.72± 1.32 0.431± 0.011 3.03± 0.22
Adaptive-OPRO -1.27± 3.21 -0.025± 0.071 6.75± 0.54 48.37± 0.10 0.466± 0.003 3.15± 0.02
Claude Sonnet 4
Baseline -4.49± 4.22 -0.134± 0.114 7.71± 1.06 13.43± 8.62 0.180± 0.121 5.52± 3.96
Reflection -3.78± 4.23 -0.115± 0.105 10.54± 1.58 5.21± 1.10 0.089± 0.026 5.11± 1.86
Adaptive-OPRO -5.07± 4.53 -0.165± 0.143 9.23± 2.71 25.85± 10.61 0.290± 0.087 3.75± 0.59
Claude Sonnet 4
w/ Thinking
Baseline -0.99± 0.80 -0.039± 0.020 7.75± 1.00 12.52± 2.47 0.175± 0.030 5.03± 1.53
Reflection -1.49± 3.76 -0.069± 0.123 7.27± 2.26 11.12± 4.86 0.186± 0.083 3.42± 2.23
Adaptive-OPRO -1.01± 0.90 -0.046± 0.020 5.16± 0.52 16.36± 7.87 0.217± 0.105 5.18± 2.52
GPT-o4-mini Baseline 1.29± 1.38 0.021± 0.044 3.23± 0.48 7.00± 3.46 0.125± 0.054 2.74± 0.79
Reflection -1.48± 0.54 -0.087± 0.018 4.64± 0.75 9.80± 3.21 0.189± 0.067 2.45± 1.00
Adaptive-OPRO 3.88± 2.21 0.089± 0.067 3.28± 0.95 10.47± 3.84 0.193± 0.046 3.42± 0.90
GPT-o3 Baseline -0.60± 1.71 -0.034± 0.050 5.93± 1.33 22.70± 0.92 0.269± 0.029 6.82± 3.03
Reflection -1.55± 2.09 -0.084± 0.075 5.02± 0.72 21.98± 4.54 0.325± 0.040 3.14± 0.99
Adaptive-OPRO 3.62± 0.90 0.096± 0.027 3.46± 0.48 25.06± 4.28 0.392± 0.019 2.31± 0.80
Table 2: Combined performance table across two markets: XOM (range-bound) and NVDA (bullish). Includes ROI, SR, and $D$. Bold values indicate the best results per model. Full results are available in Appendix E.
5 Results
Tables 1 and 2 compare three prompting strategies within ATLAS (fixed baseline, reflection, and Adaptive-OPRO) across multiple LLM families and regime-specific equity studies, alongside non-LLM baselines, under our proposed experimental design. Results show that Adaptive-OPRO consistently improves upon fixed prompts across models and market conditions, while reflection often deteriorates performance or provides inconsistent value. Non-LLM strategies demonstrate regime-dependent performance, with different technical approaches succeeding in specific conditions but failing to generalize. ATLAS with Adaptive-OPRO delivers stable performance across tested regimes, with certain model pairings achieving positive returns even in volatile and declining market conditions where most baseline strategies struggle. The order-level action space reveals distinct trading patterns across model families and supports attribution from analytical reasoning to execution behavior.
5.1 Optimization in Sequential Decision-Making
Adaptive-OPRO consistently outperforms both static baseline prompts and reflection-based approaches across the tested models and market conditions. The windowed, data-driven optimization translates into measurably better trading performance across multiple dimensions.
Return and risk-adjusted metrics demonstrate successful adaptation to market feedback. Models paired with Adaptive-OPRO achieve improved returns while maintaining or reducing drawdowns, with Sharpe ratio gains indicating that performance improvements stem from genuine strategic enhancement rather than increased risk-taking. For instance, in the volatile bearish regime (Table 1), GPT-o3 and GPT-o4-mini shift from negative baseline returns to substantial positive performance under Adaptive-OPRO, while Qwen3-235B moves from losses to positive returns. This pattern appears across diverse regimes, from volatile declining markets to range-bound and bullish conditions, suggesting that the optimization mechanism captures regime-appropriate behaviors rather than overfitting to specific conditions.
Win rate improvements reveal an impact on decision consistency. Adaptive-OPRO configurations generally achieve higher win rates alongside better returns, indicating more reliable decision-making rather than occasional large gains masking frequent losses. This consistency appears even in models that struggle under fixed prompts, suggesting that prompt evolution helps models discover and maintain effective trading patterns.
The reflection paradox emerges across models and regimes: reflection-based approaches not only fail to match Adaptive-OPRO but frequently underperform baseline prompts. This finding challenges assumptions about the universal value of reflective reasoning in sequential decision systems. When base prompts are already well-engineered, adding reflection mechanisms appears to introduce complexity without corresponding benefits, and in some cases actively degrades performance by disrupting effective decision patterns.
[…]
LLY No News 4.07± 0.72 0.056± 0.016 7.84± 3.15 53.51± 6.67 25.33± 4.51
(Bearish/Volatile No Market Data -5.75± 0.76 -0.094± 0.017 11.32± 2.63 37.52± 4.87 18.33± 3.06
Regime) No News & No Market -6.86± 1.68 -0.078± 0.036 14.54± 3.30 43.94± 6.94 22.33± 1.15
ATLAS 9.06± 0.73 0.094± 0.008 11.48± 0.00 65.28± 16.84 17.33± 5.86
XOM No News -8.20± 1.64 -0.264± 0.069 9.09± 2.99 22.82± 13.65 35.00± 12.29
(Sideways No Market Data 0.01± 0.92 -0.011± 0.021 6.56± 1.58 46.55± 23.15 13.33± 3.06
Regime) No News & No Market -4.60± 0.70 -0.136± 0.026 7.01± 2.29 35.26± 13.09 21.00± 4.58
ATLAS 3.88± 2.21 0.089± 0.067 3.28± 0.95 47.95± 7.15 25.33± 5.03
NVDA No News 6.62± 0.25 0.090± 0.008 6.67± 0.36 41.96± 5.21 28.33± 4.62
(Bullish No Market Data 11.78± 1.76 0.216± 0.024 3.70± 0.86 70.24± 14.03 20.00± 5.57
Regime) No News & No Market 7.34± 2.79 0.110± 0.012 5.76± 2.01 63.84± 9.39 20.67± 1.53
ATLAS 10.47± 3.84 0.193± 0.046 3.42± 0.90 62.70± 11.25 20.33± 2.89
Table 3: Ablation study results showing individual agent contributions using GPT-o4-mini across three market regimes. Bold values indicate the best results per configuration.
5.2 Agent Contribution Analysis
Table 3 shows distinct agent contributions through performance drops when each is ablated.
Market Analyst is a core component across market regimes. Its removal consistently results in the most significant performance degradation, especially in challenging conditions such as the bearish regime, where technical context is crucial for decision-making. In the sideways regime, the absence of market analysis not only reduces returns but also lowers trading frequency, suggesting that agents lose confidence to act without a solid technical foundation. Notably, in bullish markets, ROI slightly improves when market data are excluded, suggesting that in up-trending markets social consensus and market news may offer cleaner entry signals.
News analyst contributes regime-specific strategic value. In the bullish regime, news removal leads to lower returns as agents become more conservative, missing chances to capitalize on positive momentum. The sideways regime shows news analysis as critical, with its removal producing severe degradation-suggesting that sentiment analysis is essential when technical signals are ambiguous.
Combination of News & Market Analyst offers insights into their interdependent value: across all regimes, removing both agents leads to substantial performance degradation, indicating that news and market signals offer complementary, non-redundant information. In the bearish regime, performance drops significantly, reflecting the importance of sentiment and technical context under volatility. In the sideways regime, the absence of both leads to unstable and unprofitable behavior. Even in the bullish regime, where market data alone may be less essential, combined removal clearly harms performance. These results suggest that each component contributes differently across regimes, with the combined removal producing regime-specific effects that differ from simple additive impacts.
5.3 Trading Behavior Across LLMs
The order-level action space reveals systematic behavioral differences across model families, with performance correlating broadly with general model capabilities. Examining variance across runs provides additional insight into decision stability and reliability.
GPT-o3 demonstrates sophisticated strategy formulation, systematically integrating inputs from specialized agents into coherent trading decisions. The model exhibits conservative risk management, which limits gains in strongly trending markets but enables consistent performance across regimes. Its prompt optimization shows clear incremental learning and strategic adaptation to feedback. Notably, GPT-o3 maintains relatively low variance across runs, indicating stable decision patterns.
GPT-o4-mini favors short-term risk control through frequent stop-losses and early profit-taking. This approach proves effective in volatile conditions but struggles to capture sustained trends. The model tends toward higher trading frequencies in certain regimes, though its prompt optimization capabilities remain strong and its performance variance stays moderate.
Qwen models show divergent behavior based on scale. Qwen3-235B exhibits selective trading with moderate frequencies and achieves positive returns in multiple regimes, suggesting effective risk-reward balancing. Qwen3-32B displays higher activity levels with more variable outcomes, indicating less stable decision patterns. Both models benefit substantially from Adaptive-OPRO, with the larger variant showing particularly strong optimization responses.
LLaMA 3.3-70B adopts simpler trading strategies with limited risk management sophistication. The model shows delayed responses to market shifts and occasional abrupt strategy changes. Interestingly, this straightforward approach performs well in the bullish regime, effectively capturing upward trends without overcomplicating execution, though it struggles in more nuanced conditions.
Claude Sonnet 4 exhibits distinct behavior depending on reasoning mode. Variance patterns across runs reveal important differences in reliability: certain configurations show substantially higher standard deviations, indicating less predictable decision-making. With extended thinking enabled, the model produces detailed analysis but shows mixed translation to execution quality. Without thinking, decision patterns become even more erratic. Both configurations demonstrate challenges in maintaining consistent performance across regimes.
These behavioral patterns emerge clearly through order-level specifications, where we can separate analytical quality from execution choices. Models that struggle often show sound market analysis but poor translation into position sizing or timing, while successful configurations demonstrate coherent integration from analysis through execution. The variance patterns further distinguish models with stable, reproducible strategies from those with erratic decision-making that undermines reliability.
5.4 LLM Optimization Capabilities
A key advantage of Adaptive-OPRO is that the optimization process produces interpretable instruction updates that can be evaluated from two perspectives: whether the optimized prompt aligns with the trading objective, and whether the model successfully follows it. This dual lens reveals distinct capabilities across model families.
GPT models produce consistent, well-structured prompt refinements that demonstrate clear understanding of optimization objectives. The evolved instructions show coherent strategic adjustments that directly address observed weaknesses, and the models reliably follow their own optimized prompts.
Qwen models demonstrate strong prompt optimization capabilities across both variants. The generated refinements show analytical depth in identifying performance patterns and proposing targeted improvements, with the larger Qwen3-235B exhibiting particularly sophisticated reasoning about instruction design.
LLaMA frequently hallucinates edits, reporting modifications that do not appear in the actual prompt text or proposing changes that contradict the optimization objective. This disconnect between claimed improvements and actual instruction updates undermines the reliability of its optimization process.
Claude models tend toward overly rigid, procedural instructions that limit adaptability. The optimized prompts often become increasingly prescriptive rather than strategically focused, potentially constraining the model’s ability to respond flexibly to changing market conditions.
Reflection mechanisms, while conceptually appealing, introduce noise in practice. Even capable models can become paralyzed by over-analysis, while less sophisticated models generate vague reflections or confidently misinterpret market signals. This finding suggests that meta-reasoning about past decisions may complicate rather than clarify sequential decision-making when base prompts are already well-tuned. Further examples and analysis of these behaviors are provided in Appendix G.
6 Conclusion
We introduced ATLAS, an LLM-based trading framework integrating Adaptive-OPRO for prompt optimization under delayed, noisy feedback with structured analyst inputs and an order-level decision interface linking reasoning to execution. Across regime-specific equity studies and multiple model families, Adaptive-OPRO improves over strong fixed prompts, while reflection-based tuning yields no stable gains. The order-level interface exposes model-specific trading behaviors and separates analytical quality from execution choices, yielding clearer attribution and more interpretable decision traces. ATLAS illustrates a practical paradigm for sequential decision-making with LLMs, combining data-driven prompt evolution, structured inputs, and execution-aware outputs to achieve reliable, auditable, and accountable trading performance.
Limitations
Our scope-controlled experiments reveal how prompt optimization interacts with sequential decisions but necessarily limit breadth. We study three equities over regime-specific, two-month windows with daily decisions; sufficient to isolate adaptation effects but not to generalize across assets or market conditions. Results should thus be read as behavioral evidence of Adaptive-OPRO, not as market-wide performance claims.
Agents operate in an order-level simulator that enforces trading semantics while abstracting market microstructure. This improves control and attribution but omits execution frictions, so absolute returns may differ under richer dynamics. We use end-of-day decisions with a windowed ROI score chosen for stability rather than universality across horizons or volatility regimes.
Each configuration runs three times due to resource limits, capturing stochastic variance but limiting statistical power. Comparisons isolate adaptation under a shared interface rather than across full system architectures. Order-level actions aid interpretability and error attribution. We do not provide a causal comparison to directional-only outputs. Results span several model families (GPT, Claude, LLaMA, Qwen), but behaviors may vary with architecture, scale, or decoding choices not covered here.
Ethical Considerations
This work focuses on controlled, simulated trading experiments to study prompt optimization and does not involve real-world financial transactions or human subjects. All analyses are conducted in a reproducible, transparent environment, minimizing potential risks. While findings provide insights into model behavior, they are not financial advice and should not be used for live trading.
Appendix A Financial Markets and Trading Foundations
This appendix summarizes the trading concepts needed to interpret an order-aware interface and the signals used by the Market Analyst. The focus is on how ATLAS expresses decisions as executable orders in StockSim rather than on venue-specific microstructure.
A.1 Orders and Positions
ATLAS expresses actions at the order level and supports both long and short positioning.
Order types.
Market orders seek immediate execution at the best available prices and prioritize certainty of fill over price control. Limit orders specify a worst acceptable price for buys or a best acceptable price for sells and prioritize price control over certainty of execution. Stop orders activate once a trigger is reached and are commonly used for risk control or momentum entry.
Long and short.
A buy to open creates or increases a long position. A sell short creates a short position that profits if price declines. Exits are expressed symmetrically as sell to close for long positions and buy to cover for short positions. The Central Trading Agent may attach stops or limits to manage risk and profit-taking for either side.
Decision cadence.
The Central Trading Agent makes decisions on a daily schedule. At each decision point it consumes the updated analyst summaries and current portfolio state, then may submit new or modifying orders that are evaluated by StockSim under standard semantics. At initialization, the portfolio holds \$100,000 in cash and no positions. Since our headline metrics are percentage based (e.g., ROI, Sharpe, and drawdown computed from returns), the absolute starting capital does not affect reported performance and only scales dollar P&L.
A.2 Regime Taxonomy
We organize evaluation windows by broad market regimes in order to study behavior under distinct conditions.
Bearish volatile denotes periods with sustained downward drift and elevated variability. Sideways denotes range-bound behavior with mixed signals and limited trend persistence. Bullish denotes periods with sustained upward drift and comparatively orderly pullbacks. In the main experiments we instantiate one window for each regime and keep the decision cadence and interface fixed. The taxonomy is agnostic to any single indicator choice and can be operationalized by simple trend and volatility summaries when needed.
Appendix B Technical Indicators Used in Market Analysis
This appendix provides detailed explanations of the technical indicators employed by the Market Analyst agent in ATLAS, covering their mathematical formulations, implementation specifics, and interpretive significance in financial market analysis. All technical indicators described in this section are calculated by the StockSim Papadakis et al. (2025) simulation environment and integrated into our analysis framework to provide comprehensive market insights.
Data source.
The Market Analyst consumes OHLCV, volume, and session VWAP series from Polygon.io [^2] https://polygon.io for the specified instrument and evaluation window. Bars are retrieved at daily resolution and aligned to official U.S. market sessions, with corporate actions (splits and dividends) from Polygon used to adjust prices consistently with StockSim. All technical indicators described in this appendix are computed inside StockSim from these Polygon-derived bars. Days with incomplete or missing bars are excluded rather than backfilled, and no survivorship or lookahead adjustments are applied beyond standard split and dividend handling.
B.1 Simple Moving Average (SMA) and Exponential Moving Average (EMA)
Simple Moving Average (SMA): The SMA is calculated as the arithmetic mean of closing prices over a specified number of periods Murphy (1999):
$SMA_{n}=\frac{1}{n}\sum_{i=0}^{n-1}P_{t-i}$ (2)
where $P_{t}$ represents the closing price at time $t$ and $n$ is the number of periods. For our analysis, we employ SMA periods of 20, 50, 100, and 200 days to capture short-term, medium-term, and long-term trend characteristics. SMA provides equal weight to all prices in the calculation period, which makes it suitable for identifying longer-term trends but less responsive to recent price changes Murphy (1999).
Exponential Moving Average (EMA): The EMA assigns exponentially decreasing weights to older prices, which makes it more responsive to recent price movement Murphy (1999):
$EMA_{t}=\alpha\cdot P_{t}+(1-\alpha)\cdot EMA_{t-1}$ (3)
where $\alpha=\frac{2}{n+1}$ is the smoothing factor and $n$ is the number of periods. In our implementation, we utilize 12-period and 26-period EMAs, which serve as the foundation for MACD calculation and provide complementary trend analysis to our SMA suite. Research indicates that EMA often outperforms SMA in volatile conditions due to its enhanced sensitivity to recent price changes Kaufman (2013).
B.2 Relative Strength Index (RSI)
The RSI is a momentum oscillator that measures the speed and magnitude of price changes, oscillating between 0 and 100 Wilder (1978):
$RSI=100-\frac{100}{1+RS}$ (4)
where $RS=\frac{Average\ Gain}{Average\ Loss}$ over a specified period. Our analysis uses the standard 14-day period as originally recommended by Wilder (1978). The average gain and loss are calculated using exponential smoothing as originally formulated:
$\overline{G}_{t}=\frac{13\overline{G}_{t-1}+G_{t}}{14}$ (5)
$\overline{L}_{t}=\frac{13\overline{L}_{t-1}+L_{t}}{14}$ (6)
where $\overline{G}_{t}$ represents the average gain at time $t$, $\overline{L}_{t}$ represents the average loss at time $t$, $G_{t}$ is the current gain, and $L_{t}$ is the current loss. RSI values above 70 typically indicate overbought conditions, while values below 30 suggest oversold conditions Wilder (1978). These thresholds can be adapted to asset volatility and regime Murphy (1999).
B.3 Moving Average Convergence Divergence (MACD)
MACD is a trend-following momentum indicator that shows the relationship between two moving averages of a security’s price Murphy (1999):
$\textit{MACD}=EMA_{12}-EMA_{26}$ (7)
$\textit{Signal Line}=EMA_{9}(\textit{MACD})$ (8)
$\textit{Histogram}=\textit{MACD}-\textit{Signal Line}$ (9)
We employ the standard configuration. Crossovers and divergences are commonly used to identify trend changes and momentum shifts Achelis (2000).
B.4 Average True Range (ATR)
ATR measures market volatility by calculating the average of true ranges over a specified number of periods, as developed by Wilder (1978):
$\begin{split}\textit{True Range}=\max[(High-Low),\\ |High-Close_{prev}|,|Low-Close_{prev}|]\end{split}$ (10)
$ATR_{n}=\frac{1}{n}\sum_{i=0}^{n-1}TR_{t-i}$ (11)
We use the standard 14-period ATR. ATR supports volatility-aware sizing and stop placement.
B.5 Bollinger Bands
Bollinger Bands consist of three lines: a middle band and two outer bands positioned at standard deviations above and below the middle band Achelis (2000):
$\text{Middle Band}=SMA_{20}$ (12)
$\text{Upper Band}=SMA_{20}+(k\times\sigma)$ (13)
$\text{Lower Band}=SMA_{20}-(k\times\sigma)$ (14)
where $k$ is typically 2 and $\sigma$ is the rolling standard deviation of close. The bands adapt to changing volatility and help contextualize extremes Murphy (1999).
B.6 Support and Resistance Levels
Support and resistance levels are price zones where the asset has historically shown difficulty moving below (support) or above (resistance) Murphy (1999). We focus on horizontal levels identified by repeated interactions and elevated volume. Their strength increases with the number of tests, traded volume, and time span.
B.7 Volume Profile
Volume Profile displays trading activity over price for a chosen window:
  • Point of Control (POC): price with the highest traded volume
  • Value Area: price range that contains a specified share of volume, typically 70%
  • High Volume Nodes: locally elevated volume levels
Volume-based context helps identify zones where participation has been concentrated, which often align with support or resistance.
Appendix C Analyst Details
C.1 News Analyst
The News Analyst distills market-relevant information from financial news streams for a given ticker. Inputs are retrieved from the Polygon.io API [^3] https://polygon.io as batches of timestamped items containing title, URL, summary, and keywords. The component produces a structured analysis along four dimensions that are stable across models and assets: Sentiment Assessment, Key Developments, Market Relevance, and Source Analysis. When headline-only context is insufficient, the analyst can fetch the full article text through an internal fetcher to improve coverage and reduce headline bias. The output is designed to be compact, auditible, and directly consumable by the Central Trading Agent; it does not generate trading signals.
Example input batch (NVDA).
##NEWS BATCH
[2025-04-28T12:45:00+00:00] Want to Avoid the “Magnificent Seven” and Generate Passive Income? This Vanguard ETF May Be for You — The Motley Fool
URL: https://www.fool.com/investing/2025/04/28/magnificent-seven-passive-income-vanguard-etf/?source=iedfolrf0000001
Summary: The article discusses the Vanguard High Dividend Yield ETF (VYM) as an alternative to large-cap tech, highlighting sector diversification and dividend income.
Keywords: Vanguard High Dividend Yield ETF, Magnificent Seven, passive income, value stocks, dividend stocks
[2025-04-28T07:15:00+00:00] This Could Be a Make-It-or-Break-It Moment for Nvidia. Here’s What to Watch For This Week. — The Motley Fool
URL: https://www.fool.com/investing/2025/04/28/a-make-it-or-break-it-moment-for-nvidia/?source=iedfolrf0000001
Summary: Nvidia’s near-term outlook hinges on AI demand signals from large customers reporting earnings this week.
Keywords: Nvidia, AI, Microsoft, Amazon, Meta Platforms, Alphabet
[2025-04-27T22:25:00+00:00] Prediction: Broadcom Stock Could Surge by 100% in the Next 3 Years — The Motley Fool
URL: https://www.fool.com/investing/2025/04/27/prediction-broadcom-could-surge-by-100-in-the-next/?source=iedfolrf0000001
Summary: Broadcom’s networking and ASIC businesses are positioned to benefit from AI infrastructure demand.
Keywords: Broadcom, AI chips, networking, ASIC, VMware
Structured output (NVDA).
Sentiment Assessment Tone is cautiously optimistic and contingent on confirmation of continued AI spend by key customers. There is also a modest rotation narrative toward dividend payers that could weigh on high-beta tech.
Key Developments Earnings this week from Microsoft, Amazon, and Meta will signal near-term AI budgets. Alphabet’s reaffirmed capital spending supports sustained demand for next-gen chips. Policy risks around import tariffs remain a headwind.
Market Relevance Expect higher volatility around customer earnings. Rotation into dividend ETFs could pressure large-cap AI beneficiaries if macro uncertainty rises. Broadcom’s bullish outlook may drive sector flow dispersion that competes with Nvidia exposure.
Source Analysis All items originate from The Motley Fool, a retail-focused outlet with a constructive bias. Claims should be cross-checked against primary earnings releases and sell-side notes for actionable conviction.
Additional example (XOM).
Sentiment Assessment Mixed. ExxonMobil appears on a list of top buys for diversification strength, offset by policy uncertainty related to funding cuts for carbon capture projects.
Key Developments Federal funding for a \$332M CCS project at Baytown is being withdrawn, which may delay low-carbon hydrogen and ammonia plans, although core growth strategy remains intact.
Market Relevance Near-term noise in decarbonization headlines with limited change to base cash-flow trajectory. Integrated model and commercial partnerships support resilience.
Source Analysis Coverage from The Motley Fool blends stock-picking commentary with policy reporting and lacks direct primary citations. Verification from official releases is recommended when trading on policy moves.
Operational notes.
The News Analyst refreshes daily in sync with the decision cadence, deduplicates near-identical headlines, and preserves a consistent schema across assets and regimes. Its role is to surface catalysts, stance shifts, and source reliability in a compact form that supports downstream reasoning by the Central Trading Agent.
C.2 Fundamental Analyst
The Fundamental Analyst extracts trading-relevant structure from periodic corporate disclosures (earnings releases, financial statements) and corporate actions (dividends, splits). It runs at low frequency to mirror real reporting cadence, typically activating once or twice per evaluation window. Inputs are retrieved via Polygon.io [^4] https://polygon.io and normalized to a compact schema consumed by the Central Trading Agent. The module does not emit buy/sell signals; it summarizes material changes and likely catalysts.
C.2.1 Financial Statement Components and Terminology
Revenue and income metrics.
  • Revenue (net sales) is top-line activity prior to costs Penman (2012).
  • Gross profit margin:
    $\text{GPM}=\frac{\text{Revenue}-\text{COGS}}{\text{Revenue}}\times 100\%,$ (15)
    capturing production efficiency and pricing power Palepu et al. (2019).
  • Operating margin:
    $\text{OpM}=\frac{\text{Operating Income}}{\text{Revenue}}\times 100\%,$ (16)
    reflecting core cost discipline Penman (2012).
  • Net income is profit after all expenses, taxes, and interest.
  • Earnings per share (EPS):
    $\text{EPS}=\frac{\text{Net Income}}{\text{Weighted Avg.\ Shares}},$ (17)
    a per-share profitability anchor for valuation Damodaran (2012).
Cash-flow dynamics.
  • Operating cash flow (OCF) approximates cash generated by operations:
    $\text{OCF}=\text{NI}+\text{NCE}\pm\text{WCC},$ (18)
    where NI is net income, NCE non-cash expenses, WCC working-capital change Penman (2012).
  • Net cash flow aggregates operating, investing, and financing cash flows:
    $\text{NCF}=\text{OCF}+\text{ICF}+\text{FCF}.$ (19)
  • Capital allocation covers capex, buybacks, dividends, and debt paydown, each with distinct market implications.
Balance-sheet metrics.
  • Total assets and total equity summarize scale and residual value Palepu et al. (2019).
  • Debt-to-equity gauges leverage and risk:
    $\text{D/E}=\frac{\text{Total Debt}}{\text{Total Equity}}.$ (20)
    Higher values imply greater financial risk Damodaran (2012).
C.2.2 Corporate Actions and Structural Events
Stock splits.
Splits increase share count while proportionally reducing price (e.g., 1:2, 1:4, 1:10), often to improve perceived affordability and liquidity Baker and Powell (2012).
Dividends.
  • Cash dividends return capital to shareholders; policy signals management’s view on reinvestment vs. distribution Brealey et al. (2019).
  • Dividend yield:
    $\text{Yield}=\frac{\text{Annual Dividends Per Share}}{\text{Current Price}}\times 100\%.$ (21)
C.2.3 Analytical Dimensions
The analyst produces a concise, four-part summary focused on trading relevance: Profit & Margin Trends, Cash Flow & Capital Allocation, Balance Sheet & Leverage / Earnings Quality flags, and Catalyst Watch. Outputs are kept compact and directly auditable.
Example input batch (NVDA).
Stock Splits:
2024-06-10: 1:10  2021-07-20: 1:4  2007-09-11: 2:3  2006-04-07: 1:2
Dividends:
2025-03-12: \$0.010  2024-12-05: \$0.010  2024-09-12: \$0.010  2024-06-11: \$0.010  2024-03-05: \$0.040
Annual FY2025 (Filed: 2025-02-26):
Revenue \$130.5B; GPM 75.0%; OpM 62.4%; Net income \$72.9B; EPS \$2.94;
OCF \$64.1B; NCF \$1.3B; Assets \$111.6B; Equity \$79.3B; D/E 0.11.
Quarterly Q1 2025 (Filed: 2024-05-29):
Revenue \$26.0B; GPM 78.4%; OpM 64.9%; Net income \$14.9B; EPS \$5.98; NCF \$0.3B.
Quarterly Q2 2025 (Filed: 2024-08-28):
Revenue \$30.0B; GPM 75.1%; OpM 62.1%; Net income \$16.6B; EPS \$0.67; NCF \$1.0B.
Structured output (NVDA).
Profit & Margin Trends Q1→Q2 revenue grew $\sim$15% to \$30B on sustained AI demand; gross margin held near 75% while operating margin eased from 64.9% to 62.1%, consistent with mix normalization.
Cash Flow & Capital Allocation FY25 OCF \$64B ($\sim$49% of sales) supports heavy capex and buybacks; net cash still positive. The cut in quarterly dividend from \$0.04 to \$0.01 signals prioritization of reinvestment.
Balance Sheet & Earnings Quality Low leverage and strong equity base support flexibility. The sharp EPS swing (Q1 \$5.98 vs. Q2 \$0.67) warrants a GAAP vs. non-GAAP review to isolate one-offs.
Catalyst Watch Upcoming guidance on AI trajectory, capex cadence, and inventory dynamics are potential volatility catalysts relative to consensus.
Additional example (XOM).
Profit & Margin Trends FY2024 net margin near 10% with operating margin $\sim$14–15%; quarterly prints show stability.
Cash Flow & Capital Allocation Strong free cash flow capacity; negative annual net cash reflects investing and distribution outflows (capex, buybacks, dividends) rather than operating stress.
Balance Sheet & Leverage Debt-free posture and current ratio >>1.3 provide high financial flexibility; equity base expanded through FY/Q3.
Catalyst Watch Capital-return actions (buyback/dividend changes) and updates on large projects are the near-term fundamental triggers.
Appendix D Experiments
D.1 Evaluation Scope
We evaluate ATLAS over a two-month window (28 Apr–28 Jun 2025) across three sector-diverse equities. This horizon provides multiple decision cycles per asset while keeping full conversation histories within context limits and avoiding regime mixing. The period naturally includes routine corporate events and news, yielding a representative test bed.
D.2 Asset Selection Strategy
We use three equities chosen ex ante by simple, transparent criteria (liquidity, sector diversity, characteristic behavior): NVDA (technology, trending), LLY (healthcare, volatile drawdowns), XOM (energy, range-bound). This mix stresses different information channels and trading behaviors (trend capture, volatility management, and patience) without relying on outcome-driven selection.
D.3 Framework Configurations
Beyond the main-paper comparisons, we implemented additional variants to probe design choices:
  • Baseline: Multi-agent with carefully engineered static prompts.
  • Adaptive-OPRO: Prompt optimization applied only to the Central Trading Agent.
  • Reflection: A reviewer agent that produces periodic feedback on recent decisions. We tested weekly reflections (as in prior work) and a shorter 1-day variant; the latter is exploratory and omitted from the main tables.
  • Adaptive-OPRO + Reflection: Combined for interaction analysis; included here for completeness.
All runs keep analyst prompts fixed to isolate the adaptation mechanism at the decision layer.
D.4 Model Selection
We study how backbone capabilities translate to sequential decisions under identical interfaces:
  • Reasoning-enabled: GPT-o3, GPT-o4-mini, Claude Sonnet 4 (thinking).
  • Matched base model: Claude Sonnet 4 (no thinking) to isolate the effect of explicit reasoning.
  • Open-source: LLaMA 3.3-70B, Qwen3-235B, Qwen3-32B to gauge transfer across families and deployment options.
Within a run, the same backbone powers all ATLAS components to avoid cross-model confounds.
D.5 Ablation Study Choices
To quantify information value within ATLAS, we run ablations exclusively under GPT-o4-mini + Adaptive-OPRO:
  1. 1.
    No Market Analyst: removes multi-timescale technical structure and indicators.
  2. 2.
    No News Analyst: removes unstructured text processing of headlines and stories.
  3. 3.
    No Market & No News: leaves only portfolio state and fundamentals.
We do not ablate the Fundamental Analyst due to its intentionally low activation frequency within these windows; its role is assessed qualitatively around reporting events. Each ablation is run three times.
D.6 Evaluation Methodology
We use a multi-run protocol of three independent runs per configuration and report mean $\pm$ standard deviation. Metrics mirror the main paper (returns, risk-adjusted returns, drawdowns, win rate on closed trades, and activity). In addition to aggregate metrics, we examine decision patterns and adaptation trajectories to explain why configurations differ.
D.7 Non-LLM Based Strategies
We compare against established trading strategies (Buy & Hold, moving average crossovers, MACD) that require no machine learning. These baselines contextualize LLM performance-showing where adds value versus simpler alternatives. A detailed description of these methods is presented below.
Buy and Hold
The Buy and Hold strategy is a passive investment approach in which an asset is acquired at the beginning of the investment horizon and retained without any further trading actions, regardless of interim price fluctuations. This method assumes that, over time, the market tends to grow, and thus long-term holding can yield positive returns. It does not rely on any predictive model or technical indicator. In our evaluation, Buy and Hold serves as a benchmark strategy against which the performance of all other trading methods is compared.
Simple Moving Average (SMA)
The SMA strategy Gencay (1996) issues trading signals based on the relationship between the current price of an asset and its moving average over a fixed time window. Specifically, a buy (sell) signal is triggered when the price crosses above (below) the SMA. We test various window lengths selecting the optimal period based on validation performance.
Short-Long Moving Average (SLMA)
The SLMA method Wang and Kim (2018) extends the SMA approach by employing two SMAs of different lengths: one short-term and one long-term. A buy signal is generated when the short-term average crosses above the long-term average, while a sell signal occurs at the inverse crossover.
Moving Average Convergence Divergence (MACD)
The MACD strategy Wang and Kim (2018) captures momentum shifts by computing the difference between the 12-day and 26-day exponential moving averages. A 9-day EMA of the MACD line is used as a signal line. Trading signals are generated when the MACD line crosses the signal line from below (buy) or from above (sell). The exponential formulation ensures increased sensitivity to recent price movements.
Bollinger Bands
The Bollinger Bands strategy Day et al. (2023) incorporates volatility by constructing a band around a 20-day SMA, with the upper and lower bands placed two standard deviations above and below the mean, respectively. A price crossing above the upper band may indicate overbought conditions (sell signal), while crossing below the lower band may suggest oversold conditions (buy signal). We adopt the standard parameterization of 20-day SMA and multiplier 2, as commonly suggested in the literature.
Model Prompting Ann. SR $\uparrow$ Sortino $\uparrow$ ROIC (%) $\uparrow$ P/T (\$) $\uparrow$
LLM-Based Strategies - ATLAS
LLaMA 3.3-70B Baseline 6.16± 1.52 0.97± 0.22 30.98± 26.06 456.27± 790.29
Reflection 6.70± 0.37 1.03± 0.02 29.14± 21.06 1511.32± 2617.69
Adaptive-OPRO 6.63± 0.25 1.05± 0.01 42.26± 1.68 0.00
Claude Sonnet 4 Baseline 2.86± 1.93 0.45± 0.33 2.82± 2.60 1212.88± 920.24
Reflection 1.42± 0.41 0.16± 0.05 0.86± 0.36 416.79± 149.76
Adaptive-OPRO 4.60± 1.38 0.68± 0.22 8.25± 9.83 371.70± 1779.64
Claude Sonnet 4
w/ Thinking
Baseline 2.78± 0.48 0.46± 0.20 3.27± 1.51 1246.39± 143.77
Reflection 2.95± 1.32 0.57± 0.40 4.33± 1.72 1042.20± 424.00
Adaptive-OPRO 3.45± 1.66 0.76± 0.56 5.44± 2.81 2402.02± 1239.52
GPT-o4-mini Baseline 1.98± 0.86 0.27± 0.14 0.81± 0.39 212.27± 421.02
Reflection 3.00± 1.06 0.47± 0.23 1.40± 0.70 537.97± 45.35
Adaptive-OPRO 3.07± 0.73 0.41± 0.12 1.54± 0.47 506.75± 329.55
GPT-o3 Baseline 4.27± 0.47 0.61± 0.14 8.03± 1.86 4262.67± 897.79
Reflection 5.16± 0.63 0.68± 0.20 6.76± 2.76 2192.28± 920.54
Adaptive-OPRO 6.22± 0.30 1.22± 0.37 17.04± 7.65 3761.99± 749.07
Qwen3-235B Baseline 6.61± 0.02 0.67± 0.00 40.90± 0.35 0.00± 0.00
Reflection 5.94± 1.20 0.58± 0.14 27.90± 23.15 491.18± 850.75
Adaptive-OPRO 6.63± 0.00 0.67± 0.00 41.26± 0.00 0.00± 0.00
Qwen3-32B Baseline 7.57± 0.96 0.63± 0.07 16.37± 21.12 1567.18± 1369.31
Reflection 6.85± 0.18 0.67± 0.00 26.67± 17.99 3266.26± 5812.42
Adaptive-OPRO 7.41± 0.05 0.72± 0.01 43.27± 4.61 248.26± 200.41
Table 4: Additional performance metrics for NVDA (technology sector) comparing LLM-based approaches using ATLAS in bullish market conditions. Ann. SR = Annualized Sharpe Ratio, ROIC = Return on Invested Capital, P/T = Profit per Trade. Bold values indicate the best per model.
Model Prompting Ann. SR $\uparrow$ Sortino $\uparrow$ ROIC (%) $\uparrow$ P/T (\$) $\uparrow$
LLM-Based Strategies - ATLAS
LLaMA 3.3-70B Baseline -0.38± 0.81 -0.02± 0.06 -0.03± 0.16 -26.23± 164.36
Reflection -1.32± 0.21 -0.10± 0.01 -0.21± 0.07 -227.29± 38.58
Adaptive-OPRO -0.72± 0.19 -0.06± 0.02 -0.09± 0.03 -86.11± 31.28
Claude Sonnet 4 Baseline -2.13± 1.81 -0.17± 0.13 -0.54± 0.56 -522.11± 353.17
Reflection -1.82± 1.67 -0.14± 0.13 -0.37± 0.46 -313.67± 414.48
Adaptive-OPRO -2.62± 2.27 -0.20± 0.17 -0.80± 0.48 -576.65± 491.70
Claude Sonnet 4
w/ Thinking
Baseline -0.63± 0.32 -0.04± 0.02 -0.12± 0.10 -113.56± 89.87
Reflection -1.10± 1.94 -0.09± 0.16 -0.34± 0.85 -90.06± 311.40
Adaptive-OPRO -0.73± 0.32 -0.06± 0.02 -0.39± 0.35 -133.64± 113.58
GPT-o4-mini Baseline 0.33± 0.69 0.04± 0.08 0.16± 0.21 155.33± 202.32
Reflection -1.38± 0.29 -0.14± 0.02 -0.17± 0.05 -132.49± 87.57
Adaptive-OPRO 1.41± 1.06 0.16± 0.14 0.34± 0.26 340.47± 260.95
GPT-o3 Baseline -0.54± 0.80 -0.04± 0.07 -0.10± 0.31 -64.90± 190.96
Reflection -1.33± 1.18 -0.10± 0.08 -0.43± 0.68 -187.25± 261.18
Adaptive-OPRO 1.52± 0.43 0.15± 0.05 1.08± 0.72 380.06± 44.91
Qwen3-235B Baseline -0.70± 0.22 -0.03± 0.01 -0.43± 0.13 -437.32± 151.36
Reflection -0.59± 0.54 -0.03± 0.02 -0.34± 0.25 -334.07± 245.20
Adaptive-OPRO 0.17± 0.59 0.01± 0.03 -0.02± 0.34 -12.35± 351.54
Qwen3-32B Baseline -3.23± 0.37 -0.14± 0.02 -0.95± 0.06 -854.51± 145.41
Reflection -2.56± 0.95 -0.11± 0.04 -0.68± 0.24 -709.97± 279.41
Adaptive-OPRO -0.40± 1.14 -0.02± 0.05 0.29± 0.94 -440.76± 476.89
Table 5: Additional performance metrics for XOM (energy sector) comparing LLM-based approaches using ATLAS in stable market conditions. Ann. SR = Annualized Sharpe Ratio, ROIC = Return on Invested Capital, P/T = Profit per Trade. Bold values indicate the best per model.
Model Prompting Ann. SR $\uparrow$ Sortino $\uparrow$ ROIC (%) $\uparrow$ P/T (\$) $\uparrow$
LLM-Based Strategies - ATLAS
LLaMA 3.3-70B Baseline -1.45± 0.33 -0.09± 0.02 -1.01± 0.48 -1070.14± 634.06
Reflection -1.38± 0.39 -0.08± 0.02 -0.68± 0.20 -647.13± 141.63
Adaptive-OPRO -1.05± 0.06 -0.06 -0.47± 0.19 -472.27± 174.19
Claude Sonnet 4 Baseline -1.04± 0.48 -0.06± 0.03 -2.83± 1.13 -1920.19± 323.80
Reflection -0.91± 0.21 -0.05± 0.01 -2.66± 1.47 -1206.60± 745.08
Adaptive-OPRO 0.12± 0.28 0.01± 0.02 0.00± 0.27 -144.52± 136.78
Claude Sonnet 4
w/ Thinking
Baseline -0.68± 0.77 -0.04± 0.04 -2.65± 2.53 -2084.43± 2197.78
Reflection -1.23± 0.06 -0.08 -5.21± 1.72 -2407.54± 1345.56
Adaptive-OPRO -0.06± 0.61 -0.00± 0.04 -0.35± 0.92 -278.10± 725.32
GPT-o4-mini Baseline -0.26± 0.27 -0.02± 0.02 -0.18± 0.22 -168.13± 209.76
Reflection -0.61± 0.71 -0.04± 0.04 -0.48± 0.72 -287.24± 328.38
Adaptive-OPRO 1.49± 0.12 0.09± 0.01 1.12± 0.34 1056.49± 297.92
GPT-o3 Baseline -1.27± 0.45 -0.08± 0.02 -1.67± 1.03 -792.65± 279.17
Reflection -0.84± 0.70 -0.05± 0.04 -0.90± 0.73 -497.41± 337.21
Adaptive-OPRO 2.32± 0.76 0.16± 0.07 1.98± 0.84 799.30± 242.46
Qwen3-235B Baseline -0.09± 0.61 -0.00± 0.02 -0.23± 0.67 -495.51± 489.68
Reflection -0.78± 0.52 -0.02± 0.01 -1.41± 0.92 -1625.13± 550.55
Adaptive-OPRO 0.39± 0.31 0.01± 0.01 0.28± 0.39 66.84± 79.90
Qwen3-32B Baseline -1.39± 0.49 -0.05± 0.02 -1.01± 0.34 -1194.23± 323.67
Reflection -1.04± 0.03 -0.04± 0.01 -2.28± 2.88 -728.58± 362.80
Adaptive-OPRO -0.34± 0.34 -0.01± 0.01 -0.59± 0.37 -1213.67± 297.92
Table 6: Additional performance metrics for LLY (healthcare sector) comparing LLM-based approaches using ATLAS in volatile, declining market conditions. Ann. SR = Annualized Sharpe Ratio, ROIC = Return on Invested Capital, P/T = Profit per Trade. Bold values indicate the best per model.
Model Prompting ROI (%) $\uparrow$ Sharpe Ratio $\uparrow$ Max DD (%) $\downarrow$ Win Rate (%) $\uparrow$ Num Trades
Non-LLM-Based Strategies
Buy & Hold N/A 1.14 0.013 6.97 0.00 1
MACD N/A -0.26 -0.019 5.90 0.00 3
SMA (50-day) N/A -0.13 -0.019 5.57 0.00 3
SLMA (20/50) N/A -1.12 -0.043 5.28 0.00 2
Bollinger Bands N/A 0.00 0.000 0.00 0.00 0
LLM-Based Strategies
Llama 3.3 70B Baseline -0.42± 2.06 -0.024± 0.051 5.56± 1.08 53.48± 9.56 26.00± 2.00
Reflection -2.61± 0.77 -0.083± 0.014 6.38± 0.72 46.63± 3.15 26.33± 6.51
Adaptive-OPRO -1.10± 0.44 -0.045± 0.012 5.15± 0.71 50.00± 3.85 25.33± 1.15
Claude Sonnet 4
Baseline -4.49± 4.22 -0.134± 0.114 7.71± 1.06 37.50± 4.17 19.00± 3.46
Reflection -3.78± 4.23 -0.115± 0.105 10.54± 1.58 23.84± 8.27 18.00± 6.93
Adaptive-OPRO -5.07± 4.53 -0.165± 0.143 9.23± 2.71 31.02± 7.90 18.33± 2.52
Claude Sonnet 4
w/ Thinking
Baseline -0.99± 0.80 -0.039± 0.020 7.75± 1.00 56.28± 1.50 17.00± 5.20
Reflection -1.49± 3.76 -0.069± 0.123 7.27± 2.26 45.11± 12.6 17.00± 5.57
Adaptive-OPRO -1.01± 0.90 -0.046± 0.020 5.16± 0.52 36.2± 24.47 16.33± 2.08
GPT-o4-mini Baseline 1.29± 1.38 0.021± 0.044 3.23± 0.48 39.01± 3.61 22.67± 7.57
Reflection -1.48± 0.54 -0.087± 0.018 4.64± 0.75 32.62± 7.49 27.33± 3.06
Adaptive-OPRO 3.88± 2.21 0.089± 0.067 3.28± 0.95 47.95± 7.15 25.33± 5.03
GPT o3 Baseline -0.60± 1.71 -0.034± 0.050 5.93± 1.33 60.74± 5.59 16.33± 2.52
Reflection -1.55± 2.09 -0.084± 0.075 5.02± 0.72 42.50± 6.61 16.67± 0.58
Adaptive-OPRO 3.62± 0.90 0.096± 0.027 3.46± 0.48 71.93± 15.9 16.00± 2.65
Qwen3-235B Baseline -2.43± 0.68 -0.04± 0.01 5.72± 0.16 46.67± 5.77 11.66± 0.57
Reflection -2.02± 1.44 -0.04± 0.03 6.26± 1.77 36.51± 5.50 13.33± 2.31
Adaptive-OPRO 0.27± 1.83 0.01± 0.04 7.20± 2.09 32.86± 15.45 11± 3.61
Qwen3-32B Baseline -9.14± 1.02 -0.20± 0.02 9.82± 0.90 28.85± 17.20 21± 1.73
Reflection -7.96± 3.11 -0.16± 0.06 9.05± 2.90 40.55± 15.48 24.33± 3.05
Adaptive-OPRO -1.27± 3.21 -0.03± 0.07 6.75± 0.54 35.83± 2.57 25.67± 5.5
Table 7: Complete performance comparison between non-LLM-based and LLM-based approaches using ATLAS in range-bound market conditions (XOM, energy sector). Bold values indicate the best results per model.
Model Prompting ROI (%) $\uparrow$ SR $\uparrow$ DD (%) $\downarrow$ Win Rate (%) $\uparrow$ Num Trades
Non-LLM-Based Strategies
Buy & Hold N/A 41.30 0.409 3.16 0.00 1
MACD N/A -0.62 -0.343 0.62 0.00 1
SMA ) N/A 36.77 0.384 3.12 0.00 1
SLMA N/A 15.88 0.254 2.98 0.00 1
Bollinger Bands N/A 0.00 0.000 0.00 0.00 0
LLM-Based Strategies - ATLAS
Llama 3.3 70B Baseline 37.86± 12.31 0.388± 0.096 3.46± 0.63 20.37± 35.28 13.00± 20.78
Reflection 40.40± 1.43 0.422± 0.023 2.96± 0.34 33.33± 57.74 5.33± 6.66
Adaptive-OPRO 42.07± 1.85 0.418± 0.016 3.15± 0.02 100.00± 0.00 1.33± 0.58
Claude Sonnet 4
Baseline 13.43± 8.62 0.180± 0.121 5.52± 3.96 60.83± 12.30 21.67± 9.50
Reflection 5.21± 1.10 0.089± 0.026 5.11± 1.86 39.25± 15.79 22.33± 1.53
Adaptive-OPRO 25.85± 10.61 0.290± 0.087 3.75± 0.59 43.81± 38.37 19.00± 12.17
Claude Sonnet 4
w/ Thinking
Baseline 12.52± 2.47 0.175± 0.030 5.03± 1.53 53.30± 14.47 17.00± 2.65
Reflection 11.12± 4.86 0.186± 0.083 3.42± 2.23 77.86± 2.58 17.00± 5.00
Adaptive-OPRO 16.36± 7.87 0.217± 0.105 5.18± 2.52 68.89± 30.06 12.67± 4.04
GPT-o4-mini Baseline 7.00± 3.46 0.125± 0.054 2.74± 0.79 46.29± 3.21 18.67± 1.53
Reflection 9.80± 3.21 0.189± 0.067 2.45± 1.00 54.54± 7.92 26.33± 9.61
Adaptive-OPRO 10.47± 3.84 0.193± 0.046 3.42± 0.90 62.70± 11.25 20.33± 2.89
GPT o3 Baseline 22.70± 0.92 0.269± 0.029 6.82± 3.03 66.67± 28.87 7.33± 2.52
Reflection 21.98± 4.54 0.325± 0.040 3.14± 0.99 96.67± 5.77 18.00± 3.61
Adaptive-OPRO 25.06± 4.28 0.392± 0.019 2.31± 0.80 100.00± 0.00 9.67± 4.04
Qwen3-235B Baseline 43.91± 2.31 0.42± 0.00 3.34± 0.16 0.00± 0.00 2± 0
Reflection 34.08± 12.30 0.37± 0.08 2.98± 0.30 23.81± 41.24 11.33± 16.17
Adaptive-OPRO 41.25± 0.00 0.42± 0.00 3.16± 0.00 0.00± 0.00 2± 0
Qwen3-32B Baseline 35.75± 5.35 0.48± 0.06 2.86± 0.30 60.86± 52.71 22.33± 3.06
Reflection 41.72± 1.32 0.43± 0.01 3.03± 0.22 66.67± 57.74 10.67± 5.13
Adaptive-OPRO 48.37± 0.10 0.47± 0.00 3.15± 0.02 100.00± 0.00 18± 5
Table 8: Complete performance comparison between non-LLM-based and LLM-based approaches using ATLAS in rising market conditions (NVDA, technology sector). Bold values indicate the best per model.
Model Prompting ROI (%) $\uparrow$ SR $\uparrow$ DD (%) $\downarrow$ Win Rate (%) $\uparrow$ Num Trades
LLM-Based Strategies - ATLAS
LLaMA 3.3-70B Reflection (1d) 15.12± 9.01 0.22± 0.11 3.42± 0.70 64.88± 9.16 16± 1.73
Adaptive-OPRO w/Reflection (1d) 36.31± 6.20 0.40± 0.01 2.60± 0.92 33.33± 57.74 2± 0.58
Adaptive-OPRO 42.07± 1.85 0.42± 0.02 3.15± 0.02 100.00± 0.00 1± 0.58
Claude Sonnet 4 Reflection (1d) 6.62± 2.64 0.11± 0.06 5.14± 2.91 48.48± 2.63 15± 5.13
Adaptive-OPRO w/Reflection (1d) 24.60± 3.37 0.33± 0.05 2.39± 0.81 92.67± 7.15 17± 5.86
Adaptive-OPRO 25.85± 10.61 0.29± 0.09 3.75± 0.59 43.81± 38.37 19± 12.17
Claude Sonnet 4
w/ Thinking
Reflection (1d) 12.82± 9.97 0.21± 0.12 3.23± 2.11 50.79± 30.24 9± 2.89
Adaptive-OPRO w/Reflection (1d) 18.22± 10.21 0.23± 0.11 3.54± 0.63 53.33± 17.64 8± 2.08
Adaptive-OPRO 16.36± 7.87 0.22± 0.10 5.18± 2.52 68.89± 30.06 13± 4.04
GPT-o4-mini Reflection (1d) 3.75± 2.06 0.09± 0.03 3.24± 2.80 61.88± 11.11 30± 10.79
Adaptive-OPRO w/Reflection (1d) 4.33± 0.66 0.12± 0.02 2.36± 0.51 74.39± 2.60 30± 3.61
Adaptive-OPRO 10.47± 3.84 0.19± 0.05 3.42± 0.90 62.70± 11.25 20± 2.89
GPT-o3 Reflection (1d) 12.82± 3.94 0.25± 0.05 3.52± 1.57 82.01± 9.30 13± 2.08
Adaptive-OPRO w/Reflection (1d) 11.54± 5.63 0.24± 0.08 1.89± 0.54 73.74± 23.54 16± 4.16
Adaptive-OPRO 25.06± 4.28 0.39± 0.02 2.31± 0.80 100.00 10± 4.04
Table 9: Performance comparison of advanced prompting strategies for NVDA (technology sector) using ATLAS in bullish market conditions. Bold values indicate the best per model.
Model Prompting ROI (%) $\uparrow$ SR $\uparrow$ DD (%) $\downarrow$ Win Rate (%) $\uparrow$ Num Trades
LLM-Based Strategies - ATLAS
LLaMA 3.3-70B Reflection (1d) 0.82± 1.42 0.01± 0.02 1.62± 2.80 16.67± 28.87 8± 13.86
Adaptive-OPRO w/Reflection (1d) 0.29± 0.50 0.00± 0.00 1.96± 3.39 16.67± 28.87 12± 20.78
Adaptive-OPRO -1.10± 0.44 -0.05± 0.01 5.15± 0.71 50.00± 3.85 25± 1.15
Claude Sonnet 4 Reflection (1d) -3.76± 4.23 -0.10± 0.07 7.29± 3.08 48.81± 20.03 15± 6.08
Adaptive-OPRO w/Reflection (1d) -4.48± 3.85 -0.20± 0.16 7.16± 3.31 39.17± 20.05 14± 3.51
Adaptive-OPRO -5.07± 4.53 -0.16± 0.14 9.23± 2.71 31.02± 7.90 18± 2.52
Claude Sonnet 4
w/ Thinking
Reflection (1d) 2.40± 4.39 0.05± 0.14 4.57± 1.98 48.41± 42.35 14± 5.69
Adaptive-OPRO w/Reflection (1d) -2.84± 3.73 -0.12± 0.13 8.03± 0.89 22.62± 7.43 14± 1.53
Adaptive-OPRO -1.01± 0.90 -0.05± 0.02 5.16± 0.52 36.20± 24.47 16± 2.08
GPT-o4-mini Reflection (1d) -3.81± 2.13 -0.18± 0.06 6.54± 1.95 32.86± 8.84 38± 9.71
Adaptive-OPRO w/Reflection (1d) -1.43± 0.38 -0.09± 0.02 5.37± 3.26 41.45± 7.41 38± 5.29
Adaptive-OPRO 3.88± 2.21 0.09± 0.07 3.28± 0.95 47.95± 7.15 25± 5.03
GPT-o3 Reflection (1d) -0.97± 1.08 -0.11± 0.09 3.42± 0.58 48.21± 20.28 11± 2.65
Adaptive-OPRO w/Reflection (1d) -0.51± 0.76 -0.06± 0.03 2.71± 0.18 55.18± 16.43 17± 4.73
Adaptive-OPRO 3.62± 0.90 0.10± 0.03 3.46± 0.48 71.93± 15.99 16± 2.65
Table 10: Performance comparison of advanced prompting strategies for XOM (energy sector) using ATLAS in stable market conditions. Bold values indicate the best per model.
Model Prompting ROI (%) $\uparrow$ SR $\uparrow$ DD (%) $\downarrow$ Win Rate (%) $\uparrow$ Num Trades
LLM-Based Strategies - ATLAS
LLaMA 3.3-70B Reflection (1d) -10.59± 4.89 -0.11± 0.06 16.37± 1.97 40.47± 8.25 27± 2.65
Adaptive-OPRO w/Reflection (1d) -5.03± 0.99 -0.06± 0.02 13.18± 0.22 42.86± 7.15 26± 4.93
Adaptive-OPRO -6.16± 2.08 -0.07± 0.00 14.05± 3.33 54.36± 12.44 28± 3.21
Claude Sonnet 4 Reflection (1d) -2.98± 3.38 -0.04± 0.04 10.35± 4.47 33.33± 11.55 14± 5.20
Adaptive-OPRO w/Reflection (1d) -4.68± 4.71 -0.06± 0.06 13.07± 3.68 26.19± 8.58 15± 2.65
Adaptive-OPRO 0.35± 1.78 0.01± 0.02 14.76± 2.87 43.45± 6.27 15± 2.00
Claude Sonnet 4
w/ Thinking
Reflection (1d) -5.25± 2.34 -0.05± 0.01 15.35± 4.17 24.44± 21.43 13± 6.35
Adaptive-OPRO w/Reflection (1d) -2.07± 3.49 -0.03± 0.04 8.74± 3.77 47.62± 4.12 16± 2.52
Adaptive-OPRO -0.73± 3.82 -0.00± 0.04 12.94± 2.32 43.89± 21.11 17± 5.00
GPT-o4-mini Reflection (1d) -3.84± 2.93 -0.06± 0.04 9.61± 2.13 52.46± 2.50 32± 12.50
Adaptive-OPRO w/Reflection (1d) -1.25± 1.45 -0.04± 0.03 6.51± 2.08 41.14± 15.35 27± 3.79
Adaptive-OPRO 9.06± 0.73 0.09± 0.01 11.48 65.28± 16.84 17± 5.86
GPT-o3 Reflection (1d) 0.14± 0.56 -0.01± 0.01 6.40± 1.07 73.81± 2.06 19± 3.79
Adaptive-OPRO w/Reflection (1d) 8.05± 0.30 0.16± 0.03 4.55± 1.42 76.69± 5.03 22± 5.69
Adaptive-OPRO 9.02± 3.28 0.15± 0.05 5.33± 0.14 72.81± 17.27 20± 4.16
Table 11: Performance comparison of advanced prompting strategies for LLY (healthcare sector) using ATLAS in volatile, declining market conditions. Bold values indicate the best per model.
Figure 2: ROI across three assets using Adaptive-OPRO.
Appendix E Extended Results
This appendix consolidates additional metrics and analysis that complement the main paper’s results and experimental setup. All computations use daily portfolio returns with risk–free rate $r_{f}=0$ and are reported as mean $\pm$ standard deviation over three independent runs, consistent with the protocol described in the Experiments section.
E.1 Additional Quantitative Results
Additional Evaluation Metrics.
Beyond ROI, Sharpe Ratio, Maximum Drawdown, Win Rate, and Number of Trades, we report the following complementary measures:
Annualized Sharpe Ratio (Ann. SR):
$\text{Ann.\ SR}=\text{SR}\times\sqrt{252},$
which standardizes risk-adjusted performance to a yearly scale.
Sortino Ratio:
$\text{Sortino}=\frac{\mu-r_{f}}{\sigma_{d}},$
where $\mu$ is the mean daily return and $\sigma_{d}$ is the standard deviation of negative daily returns only. This isolates downside variability.
Return on Invested Capital (ROIC):
$\text{ROIC}=\frac{\text{Net trading profit}}{\text{Average capital deployed}}\times 100,$
which evaluates capital efficiency independent of gross exposure.
Profit per Trade (P/T):
$\text{P/T}=\frac{\text{Total net profit}}{\text{Number of trades}},$
computed on closed round trips only. This reflects average value creation per completed decision cycle and should be interpreted alongside position-level outcomes and exposure management.
E.2 Risk-Adjusted Performance Validation
Extended risk-adjusted metrics reinforce the central findings (Tables 4, 5, 6). Sortino Ratio improvements under Adaptive-OPRO indicate that gains are not driven by larger risk-taking but by better mitigation of downside variability. The effect is strongest in the bearish-volatile regime, where lower downside dispersion coincides with tighter drawdown control. ROIC consistently rises with Adaptive-OPRO across model families, showing that optimization improves the efficiency of capital deployment rather than merely increasing turnover. Improvements in P/T, when paired with higher win rates, suggest more consistent decision quality and cleaner trade selection. Since P/T excludes open positions, we interpret it jointly with exposure and drawdown metrics to avoid selection bias.
E.3 The Reflection Paradox, Revisited
Reflection mechanisms show regime- and model-dependent behavior. In multiple settings they add analysis without producing commensurate execution benefits. Across the extended metrics, reflection frequently underperforms Adaptive-OPRO and often fails to exceed strong fixed prompts. Degradations are most visible in Sortino and ROIC, where added cognitive overhead appears to introduce hesitation or inconsistent sizing. These results support the view that when base prompts and interfaces are well specified, iterative self-commentary can inject noise into otherwise coherent policies.
E.4 Architectural Performance Patterns
GPT family.
GPT-o3 exhibits the most stable risk-adjusted profile. Sortino and gains under Adaptive-OPRO align with visible drawdown compression and disciplined exposure. GPT-o4-mini benefits from optimization but shows a tendency toward over-trading in some regimes. Its risk-adjusted gains are present, yet capital efficiency can lag when trade frequency rises without proportional edge.
Qwen family.
Qwen models exhibit a scale-dependent profile. Qwen3-235B trades selectively and, under Adaptive-OPRO, achieves robust ROIC and consistent Sortino gains across regimes, especially where patience and precise timing are rewarded. Qwen3-32B is more active with higher variability; Adaptive-OPRO narrows this gap by improving risk-adjusted behavior and capital efficiency, but residual volatility in outcomes remains higher than for the larger counterpart. Reflection is particularly inconsistent for the 32B variant, where added reasoning often amplifies noise.
LLaMA 3.3-70B.
Raw returns can appear competitive in trending periods, but extended metrics reveal weaker downside control and inconsistent capital efficiency. Adaptive-OPRO reduces these gaps, yet reflection often increases variance without clear risk-adjusted gains. The pattern suggests sound high-level narrative analysis with slippage at the execution layer that optimization partially repairs.
Claude Sonnet 4 (with and without thinking).
Both modes show uneven translation from analysis to execution. With thinking enabled, the model produces detailed diagnostics, but extended metrics indicate conservative positioning that can miss trend capture, leading to modest ROIC. Without thinking, decisions are less predictable and downside risk rises. Adaptive-OPRO improves both modes but does not eliminate regime sensitivity.
E.5 Extended Prompting Strategy Analysis
Adaptation frequency effects.
Daily reflection can help in range-bound markets by encouraging restraint and tighter downside control. In trending markets it often suppresses participation, leaving upside uncaptured. Weekly reflection shows fewer short-horizon reversals but still trails Adaptive-OPRO on risk-adjusted measures (Tables 9, 10, 11).
Figure 3: Daily vs weekly reflection mechanism performance comparison across models and assets, showing ROI percentages (solid = daily, striped = weekly).
Mechanism compatibility.
Combining Adaptive-OPRO with daily reflection usually outperforms reflection alone but still underperforms pure Adaptive-OPRO. The optimization signal appears sufficient on its own, while added reflective steps introduce inconsistent edits or timing noise that dilute capital efficiency and worsen Sortino in several settings.
Summary.
Across extended metrics and regimes, Adaptive-OPRO delivers consistent improvements in downside control, capital efficiency, and per-trade value creation. Reflection provides mixed benefits and often interferes with otherwise clean optimization dynamics. Architectural differences matter: GPT-o3 and Qwen3-235B translate optimization into stable, execution-aware behavior, Qwen3-32B benefits from optimization to curb variability, LLaMA gains risk-adjusted ground but remains sensitive to execution choices, and Claude variants improve under optimization yet retain regime-dependent limitations.
Appendix F Prompt Templates
This appendix collects the verbatim prompt templates for all ATLAS agents: the Central Trading Agent (CTA), Market Analyst, News Analyst, Fundamental Analyst, the Optimizer LLM, and the Reflection Analyst. Placeholders of the form {{ variable }} are instantiated at runtime. Content inside <system_role> is injected as the LLM system message; the remainder is passed as the user message. The CTA operates on a daily decision cadence ({{ action_interval }} = 1 day). Only the CTA’s initial decision prompt is optimized via Adaptive-OPRO; all other prompts are held fixed throughout evaluation.
F.1 Central Trading Agent (CTA)
The Central Trading Agent constitutes the primary decision-making unit within the ATLAS framework, responsible for synthesizing structured analytical inputs into actionable trading directives. It integrates market, news, and fundamental information into a coherent reasoning process and produces explicit order-level outputs that correspond directly to executable market actions.
The agent’s behavior is governed by a structured prompt architecture that ensures strategic coherence while allowing adaptive responsiveness to evolving market conditions. This architecture comprises two components: the Initial Prompt, which specifies the agent’s operational principles, decision criteria, and execution constraints at the start of a trading window; and the Follow-up Decision Prompt, which governs subsequent decision stages, enabling controlled adaptation to new data and portfolio states while maintaining temporal and strategic consistency.
F.1.1 Central Agent - Initial Decision Prompt
The Initial Decision Prompt specifies the operational policy of the agent at the beginning of the trading window. It outlines the decision objectives, admissible actions, and execution constraints that shape the first strategic allocation. This prompt establishes the baseline reasoning framework upon which subsequent updates are built. The prompt is provided below.
Central Agent - Initial Prompt 1# ELITE {{ instrument }} TRADER 2**Window:** {{ window_start }} $\rightarrow$ {{ window_end }} | **Current:** {{ now }} | **Interval:** {{ action_interval }} 3 4<system_role> 5You are an elite proprietary trader running a fully-concentrated book in {{ instrument }}. 6Your goal is to maximize performance by the end of the trading window through strategic positioning. 7You are a STRATEGIC TRADER, not a day-trader. Focus on meaningful moves that align with your overall strategy. 8</system_role> 9 10## Your Toolkit & Expertise 11- Pattern recognition across multiple timeframes 12- Narrative synthesis of technical, fundamental, and sentiment inputs 13- Dynamic position sizing and risk management 14- Strategic patience and selective execution 15- Long-term performance optimization over short-term noise 16 17## Trading Philosophy 18**Strategic Patience can be your greatest ally when justified.** 19- Only act when you have high conviction and clear edge 20- Let existing positions work - avoid constant adjustments 21- Your edge comes from discipline, not frequency 22 23## Trading Toolbox 24**Order Types** 25MARKET immediate LIMIT execute at price or better STOP trigger once price crosses level 26 27**Position Actions** 28BUY open/add long SELL reduce/close long SHORT open/add short SHORT_COVER close short 29 30*(Order-type semantics follow standard brokerage definitions; interpret flexibly as conditions warrant.)* 31 32## Current Context 33{% if market_open %} 34Price: O {{ open }} H {{ high }} L {{ low }} C {{ close }} | Vol {{ volume }} 35{% else %} 36**Market Closed** orders queue for next open 37{% endif %} 38 39{% if market_analysis %}*Technical*: {{ market_analysis }}{% endif %} 40{% if news_analysis %}*News*: {{ news_analysis }}{% endif %} 41{% if fund_analysis %}*Fundamentals*: {{ fund_analysis }}{% endif %} 42{% if reflection_analysis %}*Reflection*: {{ reflection_analysis }}{% endif %} 43 44## CONSTRAINTS 45**Portfolio:** 100% concentrated in {{ instrument }} with \${{ portfolio_cash }} available cash for position sizing 46 47**Critical Rules:** 48- Never exceed available cash (\${{ portfolio_cash }}) 49- Never short more than 100% of cash balance 50- Close all short positions before {{ window_end }} 51- Unfilled orders cancel at session close - resubmit to persist 52- Decisions can be made every {{ action_interval }} 53- SELL orders are automatically limited to current long holdings - overselling is impossible 54- SHORT_COVER orders are automatically limited to current short positions - over-covering is impossible 55- System enforces position limits - you cannot accidentally create invalid positions 56 57**Portfolio Snapshot** 58Long {{ shares_long }} | Short {{ shares_short }} | Net {{ shares_net }} | Cash \${{ portfolio_cash }} 59Recent activity: {{ executed_orders }} 60 61## Decision Task 62Formulate a thesis, map key levels, gauge risk vs reward, and make your decision. 63Return either a structured order list or [] if patience best serves performance by {{ window_end }}. 64 65## Output Specification 66Return **only** a JSON array - no extra text. If no action, return []. 67[ 68 { 69 "action": "BUY | SELL | SHORT | SHORT_COVER", 70 "orderType": "MARKET | LIMIT | STOP", 71 "price": float | null, 72 "quantity": integer, 73 "explanation": "Strategic reasoning and analysis that justifies this action" 74 } 75] 76 77**CRITICAL REQUIREMENTS:** 78- EXACT values: action must be BUY|SELL|SHORT|SHORT_COVER, orderType must be MARKET|LIMIT|STOP 79- NO additional fields, NO typos, NO variations - orders will fail to place otherwise 80- Always return a JSON array (even single orders). Return [] if no action is warranted. 81- Focus on strategic positioning and end-of-window performance over tactical adjustments and noise
F.1.2 Central Agent - Follow-up Decision Prompt
The Follow-up Decision Prompt regulates the agent’s iterative reasoning process after initialization. It integrates updated analytical inputs and portfolio states to determine whether position adjustments are justified. This prompt ensures adaptive responsiveness to evolving market conditions while maintaining alignment with the initial strategic configuration. The prompt is provided below.
Central Agent - Follow-up Prompt 1# TRADING UPDATE - {{ instrument }} 2**Current:** {{ now }} 3 4Continue applying your elite trading expertise to {{ instrument }}. 5 6**Key Constraints:** 7- Never exceed cash balance (\${{ portfolio_cash }}) 8- Never short more than 100% of cash balance 9- **IMPORTANT**: Unfilled orders ALWAYS cancel at session close - resubmit to persist 10- All short positions must close before {{ window_end }} 11- SELL orders are automatically limited to current long holdings - overselling is impossible 12- SHORT_COVER orders are automatically limited to current short positions - over-covering is impossible 13 14## CURRENT CONTEXT 15**Market Data:** 16{% if market_open %} 17- Open: {{ open }} | High: {{ high }} | Low: {{ low }} | Close: {{ close }} 18- Volume: {{ volume }} 19{% else %} 20**MARKET CLOSED** 21- All outstanding orders canceled at session close 22- New orders will queue for next session open 23{% endif %} 24 25**Analyst Insights:** 26{% if market_analysis %} 27### Market Analysis 28{{ market_analysis }} 29{% endif %} 30{% if news_analysis %} 31### News Analysis 32{{ news_analysis }} 33{% endif %} 34{% if fund_analysis %} 35### Fundamentals Analysis 36{{ fund_analysis }} 37{% endif %} 38{% if reflection_analysis %} 39### Reflection Analysis 40{{ reflection_analysis }} 41{% endif %} 42 43**Portfolio Status:** 44- Long Shares: {{ shares_long }} 45- Short Shares: {{ shares_short }} 46- Net Position: {{ shares_net }} 47- Available Cash: \${{ portfolio_cash }} 48- Recent Activity: {{ executed_orders | default("None") }} 49 50## YOUR DECISION 51**Strategic Update Goal:** Decide if and how the latest developments affect your thesis and whether adjustments improve end-of-window performance. 52 53**REQUIRED JSON FORMAT:** 54[ 55 { 56 "action": "BUY|SELL|SHORT|SHORT_COVER", 57 "orderType": "MARKET|LIMIT|STOP", 58 "price": float|null, 59 "quantity": integer|null, 60 "explanation": "reasoning that synthesizes new information with your ongoing strategy" 61 } 62] 63 64**Requirements:** 65- EXACT values: action must be BUY|SELL|SHORT|SHORT_COVER, orderType must be MARKET|LIMIT|STOP 66- NO additional fields, NO typos, NO variations - orders will fail to place otherwise 67- Always return a JSON array (even single orders). If no action, return []. 68- Maintain strategic discipline while adapting to market dynamics
F.2 Market Analyst
The Market Analyst module constitutes the technical assessment layer of the ATLAS framework. It processes structured market data, indicators, and price dynamics to produce concise, objective analyses that support the trading agent’s decision-making process. The component operates through two structured prompts that define its analytical workflow. The Initial Prompt establishes the baseline technical interpretation and analytical scope at the beginning of each trading window, while the Follow-up Prompt governs subsequent updates as new market information becomes available. These prompts are presented in detail below.
F.2.1 Market Analyst - Initial Prompt
The Initial Prompt defines the baseline analytical process of the Market Analyst. It specifies the structure, scope, and format of the initial technical report, focusing on market structure, price behavior, dominant patterns, and critical levels. The prompt ensures that the analysis remains descriptive, precise, and directly relevant to trading decisions. The prompt is provided below.
Market Analyst - Initial Prompt 1# ELITE MARKET ANALYST - {{ instrument }} 2**Session:** {{ session_start }} $\rightarrow$ {{ session_end }} 3**Current:** {{ current_time }} | **Interval:** {{ action_interval }} 4 5You are an expert market analyst specializing in technical analysis. 6 7**Your analytical role:** 8- Provide objective technical analysis based on market data and indicators 9- Identify patterns, trends, and structural elements in price action 10- Present factual observations about market conditions and technical levels 11- Focus on descriptive analysis rather than predictive recommendations 12 13## MARKET DATA 14 15### Multi-Timeframe Context 16{{ extended_intervals_analysis }} 17 18### Current Session 19**OHLCV:** \${{ open_price }} / \${{ high_price }} / \${{ low_price }} / \${{ close_price }} 20**Volume:** {{ volume }} | **VWAP:** {{ vwap_str }} | **Transactions:** {{ transactions }} 21 22## TECHNICAL INDICATORS 23{{ formatted_indicators }} 24 25## YOUR ANALYSIS 26 27**Analytical Excellence Goal:** Deliver the most valuable technical insights that directly inform trading decisions. Consider what a trader most needs to know right now. 28 29**Iterative Refinement:** Think through your analysis, then refine it to ensure youre highlighting the most critical market signals and actionable price levels. Focus on what matters most for trading success. 30 31Provide analysis covering: 321. **Market Structure:** Current trend context and notable support/resistance observations 332. **Price Action:** What the current session dynamics are showing 343. **Technical Patterns:** Observable confluences and technical formations 354. **Notable Levels:** Key price levels and their technical significance 36 37**Available Technical Tools:** 38- Standard indicators: Moving averages, RSI, MACD, ATR, volume analysis 39- Advanced levels: Fibonacci retracements/extensions, pivot points, psychological levels 40- Pattern recognition: Chart patterns, candlestick formations, breakout setups 41- Volume analysis: Volume profile, VWAP deviations, volume confirmation signals 42- Consider any technical tool that helps identify actionable trading levels and signals 43 44**Response Format:** 45- Keep responses concise and direct - avoid excessive detail and repetitive explanations 46- Focus on the most critical observations only, not comprehensive analysis 47- Provide essential insights without verbose elaboration 48- Each section should be 2-3 concise sentences maximum
F.2.2 Market Analyst - Follow-up Prompt
The Follow-up Prompt manages iterative updates after the initial analysis. It enables the Market Analyst to incorporate newly available data, refresh indicator readings, and re-evaluate market conditions. This prompt maintains analytical consistency with the initial framework while highlighting only the most relevant developments for ongoing trading decisions. The prompt is provided below.
Market Analyst - Follow-up Prompt 1## MARKET UPDATE - {{ instrument }} 2**Time:** {{ current_time }} 3 4Continue your role as market analyst. Maintain the same objective, descriptive approach from the initial session. 5 6## CURRENT DATA 7**OHLCV:** \${{ open_price }} / \${{ high_price }} / \${{ low_price }} / \${{ close_price }} 8**Volume:** {{ volume }} | **VWAP:** {{ vwap_str }} | **Transactions:** {{ transactions }} 9 10## TECHNICAL INDICATORS 11{{ formatted_indicators }} 12 13**Goal:** Provide the most valuable technical insights for trading decisions. Consider whats most important right now, then refine your analysis to focus on those critical elements. 14 15Cover market structure, price action, technical setup, and key levels with emphasis on actionable insights. Keep each section to 2-3 concise sentences.
F.3 News Analyst
The News Analyst module provides the narrative and sentiment analysis layer of the ATLAS framework. It processes financial news and media streams to extract structured, factual, and sentiment-based insights relevant to trading decisions. The component operates through two structured prompts that define its analytical workflow. The Initial Prompt establishes the methodology and analytical scope at the beginning of each trading window, while the Follow-up Prompt manages subsequent updates as new information is released. These prompts are presented in detail below.
F.3.1 News Analyst - Initial Prompt
The Initial Prompt defines the baseline analytical configuration of the News Analyst. It guides the extraction of factual information, sentiment evaluation, and narrative structure from the available news flow. The prompt ensures objectivity and conciseness, focusing on actionable insights that may influence market dynamics. The prompt is provided below.
News Analyst - Initial Prompt 1# ELITE NEWS ANALYST - {{ instrument }} 2**Session:** {{ session_start }} $\rightarrow$ {{ session_end }} 3**Current:** {{ current_time }} 4 5**Your analytical role:** 6- Analyze financial news content for factual information and sentiment 7- Identify narrative trends and key developments in the news flow 8- Provide objective assessment of news relevance and credibility 9- Focus on factual analysis rather than predictive interpretations 10 11**Output Requirements:** 12- Keep responses concise and direct - avoid excessive detail and repetitive explanations 13- Focus on the most critical observations only 14- Provide essential insights without verbose elaboration 15 16**Web Search Available:** Use the web_search tool when article summaries lack detail, or you need to verify key claims. 17 18## NEWS BATCH 19{{ joined_news }} 20 21## YOUR ANALYSIS 22 23**News Intelligence Goal:** Extract the most market-relevant insights from news flow that could influence trading decisions. Consider what news elements are truly significant versus noise. 24 25**Iterative Refinement:** After analyzing the news, focus your insights on whats most actionable and relevant to current market conditions. Prioritize information that matters for trading strategy. 26 27Provide analysis focused on: 281. **Sentiment Assessment:** Whats the overall sentiment trajectory and key narrative changes? 292. **Key Developments:** What significant events or announcements are reported? 303. **Market Relevance:** How might this news content relate to market conditions? 314. **Source Analysis:** Any source reliability concerns or consensus alignment issues? 32 33**Response Format:** 34- Write in simple, direct language without jargon overuse 35- Each section should be 2-3 concise sentences maximum 36- Avoid repetitive phrasing and redundant explanations 37- No excessive formatting, bold text, or bullet point lists 38- Focus on actionable observations, not comprehensive analysis
F.3.2 News Analyst - Follow-up Prompt
The Follow-up Prompt governs iterative updates following the initial analysis. It enables the News Analyst to incorporate new articles, track evolving sentiment trends, and reassess the relevance or reliability of information sources. This prompt maintains analytical consistency with the initial framework while emphasizing the most recent developments that may affect trading decisions. The prompt is provided below.
News Analyst - Follow-up Prompt 1## NEWS UPDATE - {{ instrument }} 2**Time:** {{ current_time }} 3 4Continue your role as news analyst. Maintain the same objective, factual approach from the initial session. 5 6## LATEST NEWS BATCH 7{{ joined_news }} 8 9**Goal:** Identify the most market-moving news elements and sentiment shifts. Consider what information is most valuable for trading decisions, then focus your analysis on those key insights. 10 11Cover sentiment assessment, key developments, market relevance, and source analysis. Use web_search tool if needed for additional detail.
F.4 Fundamental Analyst
The Fundamental Analyst module provides the financial-analysis layer of ATLAS. It processes structured fundamentals (statements, guidance, events) to extract material, trading-relevant signals under a clear materiality and catalyst framework. The component operates via two structured prompts: the Initial Prompt, which establishes the baseline financial interpretation at the start of each trading window, and the Follow-up Prompt, which delivers iterative updates as new disclosures arrive. These prompts are presented below.
F.4.1 Fundamental Analyst - Initial Prompt
The Initial Prompt specifies the baseline fundamental-analysis procedure, including scope (financial health, earnings quality, balance-sheet resilience, cash-flow sustainability) and catalyst identification (events, guidance changes, corporate actions). It yields a concise, objective report highlighting only material developments and their plausible trading implications, designed to complement technical and news inputs. The prompt is provided below.
Fundamental Analyst - Initial Prompt 1# ELITE FUNDAMENTAL ANALYST - {{ instrument }} 2**Session Window:** {{ session_start }} -> {{ session_end }} 3**Current Time:** {{ current_time }} 4 5## SESSION ARCHITECTURE 6**Message Types:** 71. **Setup (this message)** - Complete framework, methodology and initial fundamentals batch 82. **Delta updates** - Compact {{ action_interval }} updates with updated fundamentals 9 10**CRITICAL:** Future deltas contain NO repeated instructions. 11All analytical frameworks must persist. 12 13You are an elite fundamental analyst with deep expertise in financial statement analysis and corporate finance. 14Your reputation is built on the ability 15to quickly identify material changes in financial health and corporate events that create trading opportunities. 16You connect the dots between financial data and market implications like a seasoned equity research professional. 17 18## ANALYTICAL PHILOSOPHY 19Your edge comes from: 20- **Financial Forensics**: Uncovering the real story behind the numbers 21- **Catalyst Recognition**: Identifying financial events that drive price action 22- **Quality Assessment**: Distinguishing between earnings quality and accounting manipulation 23- **Context Integration**: Understanding how financial health connects to market behavior 24 25## OPERATIONAL FRAMEWORK 26**Core Mission:** Extract trading-relevant insights from financial data and corporate events 27**Professional Standards:** Focus on material information that could influence trading decisions 28**Quality Approach:** Prioritize actionable insights over comprehensive analysis 29 30**Output Requirements:** 31- Keep responses concise and direct - avoid excessive detail and repetitive explanations 32- Focus on the most critical observations only 33- Provide essential insights without verbose elaboration 34 35## CURRENT FUNDAMENTALS DATA 36{{ fundamental_data }} 37 38## YOUR ANALYSIS 39 40**Response Format:** 41- Each section should be 2-3 concise sentences maximum 42- Avoid repetitive phrasing and redundant explanations 43- Focus on actionable observations, not comprehensive analysis 44 45**Fundamental Intelligence Goal:** Extract the most trading-relevant insights from financial data that could influence market decisions. Consider which fundamental factors are most likely to impact price action in the current market environment. 46 47**Iterative Analysis:** Review the financial data thoroughly, then focus your insights on the most material changes and catalysts. Prioritize information that provides valuable context for trading strategy. 48 49Apply your fundamental analysis expertise to extract trading-relevant insights. Focus on corporate events, financial health trends, and performance indicators that could influence short-term trading decisions. 50 51Consider earnings quality, balance sheet strength, cash flow sustainability, and any material changes that could serve as catalysts. Your analysis should provide fundamental context that complements technical and sentiment analysis. 52 53**Remember:** Identify fundamental factors that could influence price action. Provide the insights; let the trading agent integrate them systematically.
F.4.2 Fundamental Analyst - Follow-up Prompt
The Follow-up Prompt governs incremental updates after initialization. It incorporates newly released fundamentals (filings, guidance, event deltas), reassesses material changes and catalysts, and refines the prior assessment while preserving methodological consistency. Emphasis is placed on short-horizon relevance and actionable context for the trading agent. The prompt is provided below.
Fundamental Analyst - Follow-up Prompt 1## FUNDAMENTAL ANALYSIS UPDATE - {{ instrument }} 2**Timestamp:** {{ current_time }} 3 4Continue with your role as elite fundamental analyst. Apply the same analytical depth and professional standards established in the initial framework. 5 6## UPDATED FUNDAMENTALS 7{{ fundamental_data }} 8 9**Goal:** Identify the most significant fundamental developments and their potential market implications. Consider what fundamental information is most valuable for current trading context, then focus your analysis accordingly. 10 11Provide fundamental analysis focusing on material changes and trading implications.
F.5 Trading Prompt Optimizer (Adaptive-OPRO Target = CTA Initial Prompt)
The Trading Prompt Optimizer is the meta-policy that revises only the static instruction block of the Central Trading Agent’s Initial Decision Prompt. At each window boundary it consumes a prompt–performance history (history_text) scored via the windowed ROI signal and proposes an edited template that preserves all placeholders ({{...}}), conditional blocks ({% if %}), and the order JSON schema (actions and order types). The optimizer returns a strictly structured JSON payload containing a diagnostic performance_analysis, a full optimized_prompt (template text, not a filled instance), key_improvements, and an expected_impact. An update is applied only if the placeholder set and interface remain unchanged, ensuring compatibility with the runtime injector.
[…]
F.6 Weekly Reflection Agent
The Weekly Reflection Agent provides periodic ({{reflection_interval}}-day) reviews of recent trades and portfolio evolution, producing a single, compact paragraph that highlights recurring patterns, risk discipline, and thesis maintenance. Its output is advisory text only: it is injected as reflection_analysis for the Central Trading Agent to read on subsequent decisions, and it does not directly edit prompts or alter execution semantics. The reflection is derived from the full decision log and period summary, avoids prescriptive rules or rigid thresholds, and is designed to surface durable process improvements rather than post-hoc trade-by-trade commentary. By construction, it respects the fixed decision interval and order-cancellation rules described in the environment specification.
Weekly Reflection Agent’s Prompt 1# ELITE TRADING COACH - {{ instrument }} INTERVAL REVIEW 2**Period:** {{ reflection_interval }}-day review | **Session:** {{ current_time }} | **Trading Decision Frequency:** {{ action_interval }} 3 4You are a reflection agent analyzing {{ reflection_interval }} days of trading performance to provide strategic insights for systematic improvement. 5 6## TRADING SYSTEM RULES & LIMITATIONS 7**Portfolio & Operational Context:** 8**Single-Stock Portfolio:** The agent manages a concentrated portfolio dedicated exclusively to {{ instrument }} - all available capital and positions are focused on this one security with no diversification across multiple stocks. 9**Available Actions:** BUY, SELL, SHORT, SHORT_COVER 10**Order Types:** MARKET, LIMIT, STOP 11**Constraints:** Cash limits, position sizing rules, and {{ action_interval }} decision intervals apply 12**Position Limits:** SELL orders are automatically limited to current holdings, and SHORT_COVER orders are automatically limited to current short positions - overselling or over-covering is impossible. The system enforces these limits automatically. 13**Critical Constraint:** The agent can only make trading decisions at fixed {{ action_interval }} intervals. All orders in the decision JSON are placed simultaneously - there is no sequential order placement. 14**Order Auto-Cancellation:** Unfilled orders are automatically cancelled at the end of each decision interval. 15 16## PERIOD PERFORMANCE OVERVIEW 17{{ period_summary }} 18 19## COMPLETE DECISION HISTORY FOR PERIOD 20{{ complete_history }} 21 22## YOUR COACHING TASK 23 24PURPOSE 25In one comprehensive paragraph, synthesize the most impactful patterns from this {{ reflection_interval }}-day period and identify the single structural improvement that would most enhance future performance cycles. 26Focus on systematic insights that will compound over multiple {{ reflection_interval }}-day periods rather than individual trade critiques. 27 28GUIDELINES 29- Analyze decision patterns, risk management consistency, and strategic evolution across the period 30- Identify the highest-leverage behavioral or strategic adjustment for future periods 31- Emphasize enduring principles over isolated performance details 32- Skip grades, personality assessments, or motivational language 33 34**REQUIRED OUTPUT FORMAT:** Return only your reflection as a single paragraph of continuous plain text (3-5 sentences).
Appendix G Prompt Evolution Mechanism Analysis
The transparent optimization traces produced by Adaptive-OPRO provide unprecedented insights into how systematic prompt refinement drives performance improvements in sequential decision-making systems. Through detailed examination of optimization trajectories across different model architectures, we can observe the precise mechanisms by which prompt modifications translate into enhanced trading performance.
G.1 Systematic Weakness Detection and Resolution
The optimization process demonstrates sophisticated analytical capabilities in identifying prompt weaknesses and prescribing targeted improvements. Analysis of the GPT-o3 optimization trajectory from iteration 4 to iteration 5 on LLY stock reveals the systematic approach employed by the meta-optimization process.
G.1.1 Phase 1: Diagnostic Analysis - Identifying Performance Bottlenecks
Performance Analysis: Weakness Detection Optimizer’s Weakness Identification: “Across iterations, performance rose from 43.2 → 56.6 as prompts became more concise, structured, and decision-oriented. Gains came from: (1) cleaner sectioning that reduced cognitive load, (2) explicit reasoning frameworks that guided probability-weighted thinking, and (3) clearer constraint reminders that prevented rule breaches. Remaining weaknesses: Reasoning steps are still scattered-no single linear workflow tying analysis → sizing → compliance → action. Risk-management is mentioned but not enforced with a final checklist, so occasional oversizing or sub-optimal reward-to-risk trades slip through. The JSON spec is sometimes buried deep in the prompt; occasional format errors could still occur. Context blocks (technical, news, fundamentals, reflection) are informative but not explicitly referenced in the reasoning flow, so the model may overlook one dimension.”
The optimizer’s analysis demonstrates pattern recognition across multiple iterations, identifying four critical areas for refinement: workflow linearization to create more structured reasoning chains, risk management formalization to enforce disciplined decision-making, output specification prominence to reduce formatting errors, and context integration enhancement to ensure comprehensive information utilization. This diagnostic precision enables targeted remediation rather than broad, inefficient modifications.
G.1.2 Phase 2: Strategic Intervention - Translating Insights into Targeted Solutions
Building directly upon these identified weaknesses, the optimization process prescribes specific structural modifications designed to address each diagnostic finding systematically:
Key Improvements: Targeted Solutions Strategic Modifications Implemented: 1. Introduced a 5-step THINK $\rightarrow$ CHECK $\rightarrow$ ACT workflow that linearly converts market inputs into compliant orders, minimizing reasoning omissions. 2. Added an explicit PRE-ORDER RISK CHECKLIST (cash, short limit, catalyst validity, $\geq$ 2:1 R:R) to curb rule violations and low-edge trades. 3. Elevated the four context feeds (technical, news, fundamentals, reflection) into a single MARKET SITUATION dashboard that the workflow must reference, ensuring holistic analysis. 4. Moved the strict JSON schema into its own boxed section immediately before output instructions; this reduces formatting errors. 5. Kept language concise but directive, reinforcing trader autonomy while preventing over-trading with a PATIENCE override. 6. Preserved every required {{placeholder}} and {% if %} block exactly, guaranteeing template compatibility.
Each modification directly corresponds to a specific weakness identified in the diagnostic phase, creating a clear causal chain from problem identification to solution implementation. The architectural changes shown in Figures 4, 5, and 6 demonstrate this systematic approach, consolidating scattered elements while strengthening decision-making frameworks.
G.1.3 Phase 3: Outcome Assessment - Connecting Solutions to Impact
Having implemented these targeted architectural improvements, the optimization process generates forward-looking performance predictions based on the expected behavioral changes from each modification:
Expected Impact: Performance Prediction Forward-Looking Impact Assessment: “The linear THINK $\rightarrow$ CHECK $\rightarrow$ ACT workflow anchors the model’s reasoning, reducing skipped steps and improving decision quality. The explicit risk checklist enforces discipline, likely lowering drawdowns and boosting risk-adjusted returns. Consolidating all market feeds into one dashboard ensures holistic analysis, while the clearer JSON spec lowers formatting errors. Collectively, these improvements should enhance comprehension, deepen analysis, and translate into higher-scoring, more profitable trading decisions.”
This prediction proves accurate, as performance improved from 56.6 to 67.6 following these modifications, validating the optimizer’s analytical capabilities and demonstrating the effectiveness of systematic architectural refinement.
Header and Trader Identity Evolution (Prompt 4 to Prompt 5) 1- # {{ instrument }} ALPHA COMMAND CENTER 2+ # {{ instrument }} ALPHA STRATEGY HUB 3**Window:** {{ window_start }} $\rightarrow$ {{ window_end }} | **Current:** {{ now }} | **Interval:** {{ action_interval }} 4Your singular -objective +mission is to maximise risk-adjusted performance 5by {{ window_end }} through disciplined, high-conviction positioning. Balance strategic patience with decisive execution; ignore noise. 6 7============================== 8- 1. MISSION 9+ 1. MISSION & KPI 10============================== 11Deliver superior returns while preserving capital (+by {{ window_end }}). 12- Act only when probability and reward justify the risk. 13+ Success metric: cumulative risk-adjusted performance. 14 15============================== 16- 2. YOUR EDGE 17+ 2. EDGE & PRINCIPLES 18============================== 19 Multi-timeframe pattern recognition 20 Integration of technical, fundamental & sentiment narratives 21 Dynamic risk management and position sizing 22- Capacity to remain inactive until odds are favourable 23+ Patience until odds are clearly favourable
Figure 4: Header and trader identity modifications between iteration 4 and iteration 5, showing title changes and mission statement refinements. Lines in red with a leading “-” and lines in green with a leading “+” indicate deletions and additions, respectively, proposed by Adaptive-OPRO.
Information Architecture and Constraints Consolidation (Prompt 4 to Prompt 5) 1- 3. MARKET DASHBOARD 2+ 3. MARKET SITUATION DASHBOARD 3============================== 4{% if market_open %} Price: O {{ open }} H {{ high }} L {{ low }} C {{ close }} | Vol {{ volume }}{% else %} **Market Closed** - orders queue for next open {% endif %} 5{% if market_analysis %}*Technical*: {{ market_analysis }}{% endif %} 6{% if news_analysis %}*News*: {{ news_analysis }}{% endif %} 7{% if fund_analysis %}*Fundamentals*: {{ fund_analysis }}{% endif %} 8{% if reflection_analysis %}*Reflection*: {{ reflection_analysis }}{% endif %} 9 10============================== 11- 4. OPERATING CONSTRAINTS 12- ============================== 13- Portfolio cash: \${{ portfolio_cash }} | Concentrated in {{ instrument }} only 14- Never exceed available cash 15- May short up to 100% of cash (must be flat by {{ window_end }}) 16- Unfilled orders cancel at session close 17- Decision frequency: every {{ action_interval }} 18- System blocks quantities beyond current exposure (cannot oversell or over-cover) 19 20- ============================== 21- 5. PORTFOLIO SNAPSHOT 22+ 4. PORTFOLIO & CONSTRAINTS 23============================== 24Long {{ shares_long }} | Short {{ shares_short }} | Net {{ shares_net }} | Cash \${{ portfolio_cash }} 25Recent activity: {{ executed_orders }} 26+ Never exceed available cash (\${{ portfolio_cash }}) 27+ May short up to 100% of cash (flat by {{ window_end }}) 28+ Unfilled orders cancel at session close 29+ Decision cadence: every {{ action_interval }} 30+ System blocks invalid quantities (cannot oversell/over-cover)
Figure 5: Structural reorganization consolidating sections into a unified PORTFOLIO & CONSTRAINTS section. Lines in red with a leading “-” and lines in green with a leading “+” indicate deletions and additions, respectively, proposed by Adaptive-OPRO.
Workflow Restructuring and Output Specification Enhancement (Prompt 4 to Prompt 5) 1- 6. DECISION PROTOCOL 2+ 5. THINK $\rightarrow$ CHECK $\rightarrow$ ACT WORKFLOW 3============================== 4- REVIEW $\rightarrow$ REASON $\rightarrow$ RESPOND 5- 1. REVIEW: Regime, key drivers, levels, catalysts. 6- 2. REASON: Probability map, $\geq$2:1 reward-to-risk, position sizing within constraints. 7- 3. RISK CHECKLIST: (a) Exposure aligns with conviction; (b) Catalyst still valid; (c) Downside defined & acceptable. 8- 4. RESPOND: ACT (issue order) or WAIT/HOLD. Patience is edge when conditions are unclear. 9+ STEP 1: Diagnose Regime & Narrative (use all dashboard feeds). 10+ STEP 2: Map Key Levels & Catalysts; assign probabilities. 11+ STEP 3: Define Reward:Risk (target $\geq$2:1) and provisional size within constraints. 12+ STEP 4: PRE-ORDER RISK CHECKLIST 13+ Cash / short limits respected 14+ Position aligns with conviction & catalyst 15+ Downside defined; R:R $\geq$2:1 16+ Flat by {{ window_end }} if short 17+ STEP 5: DECIDE 18+ ACT: issue orders 19+ WAIT/HOLD: output [] (patience override) 20 21============================== 22- ORDER OUTPUT SCHEMA (STRICT) 23+ 6. ORDER OUTPUT SPEC (STRICT) 24============================== 25Return ONLY a JSON array or [] - no extra text. 26Each object must match exactly: 27{ 28 "action": "BUY | SELL | SHORT | SHORT_COVER | ", 29 "orderType": "MARKET | LIMIT | STOP", 30 "price": float | null, 31 "quantity": integer, 32 "explanation": "Brief strategic reasoning" 33} 34Invalid fields, casing, or additional text will cause order rejection.
Figure 6: Decision protocol restructuring from informal REVIEW → REASON → RESPOND to structured five-step THINK → CHECK → ACT workflow. Lines in red with a leading “-” and lines in green with a leading “+” indicate deletions and additions, respectively, proposed by Adaptive-OPRO.
G.2 Progressive Prompt Evolution: From Generic Foundation to Optimized Performance
The GPT-o4-mini optimization trajectory demonstrates systematic prompt evolution through three distinct phases, each building upon previous discoveries to achieve cumulative performance improvements. The optimization process adapts to both model-specific response patterns and varying market regime requirements.
The progression from baseline (37.2) through intermediate optimization (51.4) to final optimization (72.1) reveals how systematic refinement can compound initial improvements into substantial performance gains. These three representative prompts (Prompt 1, Prompt 4, and Prompt 11) from the full optimization trajectory illustrate the key evolutionary patterns that drive performance enhancement.
The baseline prompt (Prompt 1) is documented Appendix F; here we present only the intermediate and final optimized variants to avoid duplication.
The intermediate optimization achieves structural refinement by systematically eliminating architectural complexity while strengthening core functionality. Figure 7 reveals this transformation: verbose explanations are stripped away and replaced with a compact, numbered decision framework that provides clear analytical guidance. The constraint presentation undergoes similar streamlining, retaining comprehensive coverage while dramatically improving clarity. Crucially, the framework maintains an advisory approach (Define thesis & edge) that guides without constraining, avoiding over-specification that could limit model flexibility. This architectural simplification creates a foundation optimized for further enhancement.
The final optimization achieves breakthrough performance by expanding upon this concise foundation with granular procedural guidance. Figure 8 showcases the evolved architecture where the decision framework expands to six numbered steps with explicit descriptions: Define Thesis & Edge: state your core conviction and Validate Compliance: ensure all constraints are met before submission. The market context integration becomes systematically organized with consistent bullet-point formatting and descriptive labels like Technical Analysis and News Impact. The constraint presentation achieves optimal balance between completeness and clarity, providing comprehensive operational guidance without cognitive overload. This final optimization demonstrates how systematic refinement can compound architectural improvements into substantial performance gains, with each evolution building upon and enhancing previous discoveries.
Intermediate Optimization (Prompt 4) - Score: 51.4
# ELITE {{ instrument }} TRADER
Window: {{ window_start }} → {{ window_end }} | Current: {{ now }} | Interval: {{ action_interval }}
You are an elite proprietary trader managing a fully-concentrated book in {{ instrument }}.
Maximize end-of-window risk-adjusted performance through high-conviction, strategic trades
with disciplined execution and strategic patience.
## MARKET CONTEXT
{% if market_open %}
Price: O {{ open }} H {{ high }} L {{ low }} C {{ close }} | Vol {{ volume }}
{% else %}
Market Closed - orders queue for next session
{% endif %}
{% if market_analysis %}* Technical: {{ market_analysis }}{% endif %}
{% if news_analysis %}* News: {{ news_analysis }}{% endif %}
{% if fund_analysis %}* Fundamentals: {{ fund_analysis }}{% endif %}
{% if reflection_analysis %}* Reflection: {{ reflection_analysis }}{% endif %}
## DECISION FRAMEWORK
1. Define thesis & edge
2. Identify entry, stop, and target levels
3. Assess risk/reward & size within cash limits
4. Choose order type & execution timing
5. Verify constraints & finalize plan
## CONSTRAINTS & PORTFOLIO
- Fully concentrated in {{ instrument }}, Cash \${{ portfolio_cash }}
- Long {{ shares_long }} | Short {{ shares_short }} | Net {{ shares_net }}
- Recent orders: {{ executed_orders }}
- Max short = 100% cash; close all shorts by {{ window_end }}
- Actions: BUY, SELL, SHORT, SHORT_COVER
- Order Types: MARKET, LIMIT, STOP
- Orders expire at session close; re-submit if needed
## OUTPUT SPECIFICATION
Return only a JSON array of orders or []:
    "action": "BUY | SELL | SHORT | SHORT_COVER",
    "orderType": "MARKET | LIMIT | STOP",
    "price": float | null,
    "quantity": integer,
    "explanation": "Concise strategic reasoning"
Figure 7: Intermediate optimization (GPT-o4-mini, Prompt 4) featuring streamlined structure with a numbered decision framework and concise constraint presentation. Score: 51.4
Optimized Prompt (Prompt 11) - Score: 72.1
# ELITE {{ instrument }} TRADER
Window: {{ window_start }} → {{ window_end }} | Now: {{ now }} | Interval: {{ action_interval }}
You are an elite proprietary trader managing a fully concentrated book in {{ instrument }}.
Your mission is to maximize end-of-window risk-adjusted performance with high conviction
and disciplined execution. Think critically and adapt your approach as market conditions evolve.
## MARKET CONTEXT
{% if market_open %}
- Price: O {{ open }} H {{ high }} L {{ low }} C {{ close }} | Vol {{ volume }}
{% else %}
- Market Closed - orders queue for next session
{% endif %}
{% if market_analysis %}- Technical Analysis: {{ market_analysis }}{% endif %}
{% if news_analysis %}- News Impact: {{ news_analysis }}{% endif %}
{% if fund_analysis %}- Fundamental Overview: {{ fund_analysis }}{% endif %}
{% if reflection_analysis %}- Reflection: {{ reflection_analysis }}{% endif %}
## PORTFOLIO & CONSTRAINTS
- Total Allocation: 100% in {{ instrument }}, Cash \${{ portfolio_cash }}
- Positions: Long {{ shares_long }}, Short {{ shares_short }}, Net {{ shares_net }}
- Recent Activity: {{ executed_orders }}
- Max short = 100% cash; all shorts must close by {{ window_end }}
- Orders expire at session close; unfilled orders cancel (re-submit to persist)
## DECISION FRAMEWORK
1. Define Thesis & Edge: state your core conviction.
2. Map Key Levels: identify entry, stop-loss, and target levels.
3. Assess Risk/Reward: compute per-share risk, total risk, and reward potential.
4. Allocate Size: determine quantity within cash limits (\${{ portfolio_cash }}).
5. Choose Execution: select action (BUY | SELL | SHORT | SHORT_COVER)
   and orderType (MARKET | LIMIT | STOP).
6. Validate Compliance: ensure all constraints are met before submission.
## OUTPUT SPECIFICATION
Return only a JSON array of orders or an empty array ([]). No extra text:
    "action": "BUY | SELL | SHORT | SHORT_COVER",
    "orderType": "MARKET | LIMIT | STOP",
    "price": float | null,
    "quantity": integer,
    "explanation": "Concise strategic reasoning"
Figure 8: Final optimized prompt (GPT-o4-mini, Prompt 11) with a six-step decision framework and systematic market context organization. Score: 72.1
Appendix H Reproducibility
All experiments are conducted on a MacBook Pro with an Apple M3 Pro chip (11-core CPU) and 18 GB of unified memory. Our experiments are conducted using an updated version of the StockSim environment Papadakis et al. (2025), with modifications to support the ATLAS multi-agent architecture, Adaptive-OPRO optimization, and reflection-based mechanisms (implementation details in code). An example configuration for GPT-o4-mini using Adaptive-OPRO on XOM is provided under configs/o4-mini-adaptive-opro-config.yaml. All other experimental configurations can be reproduced by following the StockSim documentation and adapting this sample.
Model ID Model Card / Provider Identifier
LLaMA 3.3-70B meta.llama3-3-70b-instruct-v1:0
Claude Sonnet 4 anthropic.claude-sonnet-4-20250514-v1:0
Qwen3 235B A22B 2507 qwen.qwen3-235b-a22b-2507-v1:0
Qwen3 32B (dense) qwen.qwen3-32b-v1:0
Table 12: Models accessed via Amazon Bedrock.
Model ID Model Card / Docs
GPT-o4-mini gpt-4o-mini-2024-07-18
GPT-o3 gpt-o3-2025-04-16
Table 13: Models accessed via OpenAI.
We access LLaMA, Claude, and Qwen models via Amazon Bedrock (Table 12). GPT models are accessed via OpenAI APIs (Table 13). We interface with all LLMs strictly through provider APIs and do not employ any local hardware or fine-tuning.